### v1

# Computational bottlenecks for denoising diffusions

Abstract
Denoising diffusions provide a general strategy to sample from a probability distribution $\mu$ in $\mathbb{R}^{d}$ by constructing a stochastic process $(\hat{\boldsymbol{x}}_{t}:t\geq 0)$ in ${\mathbb{R}}^{d}$ such that the distribution of $\hat{\boldsymbol{x}}_{T}$ at large times $T$ approximates $\mu$. The drift $\boldsymbol{m}:{\mathbb{R}}^{d}\times{\mathbb{R}}\to{\mathbb{R}}^{d}$ of this diffusion process is learned from data (samples from $\mu$) by minimizing the so-called score-matching objective. In order for the generating process to be efficient, it must be possible to evaluate (an approximation of) $\boldsymbol{m}(\boldsymbol{y},t)$ in polynomial time.
Is every probability distribution $\mu$, for which sampling is tractable, also amenable to sampling via diffusions? We provide evidence to the contrary by constructing a probability distribution $\mu$ for which sampling is easy, but the drift of the diffusion process is intractable —under a popular conjecture on information-computation gaps in statistical estimation. We further show that any polynomial-time computable drift can be modified in a way that changes minimally the score matching objective and yet results in incorrect sampling.
1 Introduction
1.1 Background
Denoising diffusions [SE19, HJA20] have emerged as a central paradigm in generative artificial intelligence (AI). Given a target distribution $\mu$ on ${\mathbb{R}}^{d}$, we want to sample $\boldsymbol{x}\sim\mu$. Diffusions achieve this goal by generating trajectories of a stochastic process $(\hat{\boldsymbol{x}}_{t}:0\leq t\leq T)$ whose state $\hat{\boldsymbol{x}}_{T}$ at large $T$ is approximately distributed according to $\mu$. Crucially, the process $(\hat{\boldsymbol{x}}_{t}:0\leq t\leq T)$ is a diffusion whose coefficients can be estimated —at least in principle— from data $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{N}\sim\mu$ by solving a regression problem (empirical risk minimization with respect to square loss). Explicit knowledge of $\mu$ is not needed in order to generate a fresh sample $\boldsymbol{x}\sim\mu$ and powerful deep learning models can be deployed to approximate the coefficients of the diffusion. Even this brief description suggests a natural question:
Are there probability distributions $\mu$ for which generating the denoising diffusion is intrinsically harder than only generating samples from $\mu$?
This paper points at a positive answer. We construct a simple family of distributions $\mu$ for which: $(1)$ Sampling from $\mu$ is easy; $(2)$ Evaluating the coefficients of the diffusion is expected to be hard. More precisely, evaluation of these coefficients requires solving a statistical estimation problem for which a gap is expected to exist between optimal estimation error, and error achieved by polynomial time algorithms. We further show that any drift that is is nearly optima among polynomial time computable ones, can be modified in a way that leaves the score-matching objective nearly unchanged but results in incorrect sampling. Our construction is novel in this context and formalizes past observations based on statistical physics; see [MRTS07, RTS09, GDKZ24, HMP24] and Section 2.
We emphasize that this failure of denoising diffusions is computational of nature and is not driven by the difficulty of estimating the coefficients of the diffusion.
Denoising diffusions [^1] We follow the construction given in [Mon23], rather than the original one: the two differ by a change of variables. implement a discretization of the following stochastic differential equation (SDE), with initialization $\boldsymbol{y}_{0}=\boldsymbol{0}$
${\rm d}{\boldsymbol{y}_{t}}$ $=\boldsymbol{m}(\boldsymbol{y}_{t};t){\rm d}{t}+{\rm d}{\boldsymbol{B}_{t}},$ (1)
$\boldsymbol{m}(\boldsymbol{y},t)$ $:=\mathbb{E}\{\boldsymbol{x}|t\boldsymbol{x}+\sqrt{t}\boldsymbol{g}=\boldsymbol{y}\}\,,$ (2)
where $(\boldsymbol{B}_{t})_{t\geq 0}$ is a standard Brownian motion and expectation in the second line is with respect to $\boldsymbol{x}\sim\mu$ independent of $\boldsymbol{g}\sim{\mathsf{N}}(\boldsymbol{0},\boldsymbol{I}_{d})$.
It is not hard to show that, if $\boldsymbol{y}_{t}$ is generated according the the above SDE, then there exists $\boldsymbol{x}\sim\mu$ and an independent standard Brownian motion $(\boldsymbol{W}_{t})_{t\geq 0}$ (different from $(\boldsymbol{B}_{t})_{t\geq 0}$) such that
$\boldsymbol{y}_{t}=t\,\boldsymbol{x}+\boldsymbol{W}_{t}\,.$ (3)
Therefore, running the diffusion (1) until some large time $T$, and returning $\boldsymbol{y}_{T}/T$ or $\boldsymbol{m}(\boldsymbol{y}_{T},T)$ yields a sample approximately distributed according to $\mu$.
A standard Euler discretization of the SDE (1) is given by (for $\hat{\boldsymbol{z}}_{t}\sim{\mathsf{N}}(\boldsymbol{0},\boldsymbol{I}_{d})$ independent of the past):
$\hat{\boldsymbol{y}}_{t+\Delta}$ $=\hat{\boldsymbol{y}}_{t}+\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{t},t)\Delta+\sqrt{\Delta}\,\hat{\boldsymbol{z}}_{t}\,,$ (4)
with $\Delta$ a small stepsize. The algorithm outputs $\hat{\boldsymbol{x}}_{T}=\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{T},T)$ for some suitably large $T$. Here $\hat{\boldsymbol{m}}(\,\cdot\,,t)$ is the estimated drift. We will refer to this as ‘diffusion sampling.’
Diffusions reduces the problem of sampling to the problem of approximating the conditional expectation (2). The mapping $\boldsymbol{y}\mapsto\boldsymbol{m}(\boldsymbol{y},t)$ is the Bayes-optimal estimator of $\boldsymbol{x}$ in Gaussian noise:
$\boldsymbol{m}(\,\cdot\,,t)=\underset{{\boldsymbol{\varphi}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}}{\operatorname{argmin}}\;\mathbb{E}\big{\{}\|{\boldsymbol{\varphi}}(\boldsymbol{y}_{t})-\boldsymbol{x}\|^{2}\big{\}}\,.$ (5)
In words, we are given a Gaussian observation $\boldsymbol{y}_{t}\sim{\mathsf{N}}(t\boldsymbol{x},t{\boldsymbol{I}}_{d})$ (for a single $t$) and want to estimate $\boldsymbol{x}$ as to minimize mean square error (MSE). Here the minimization is over all measurable functions ${\boldsymbol{\varphi}}$.
In practice we are given a finite sample $(\boldsymbol{x}_{i})_{i\leq N}\sim_{iid}\mu$. We thus replace the expected error empirical error and the minimization over all functions, by minimization over a suitable function class:
minimize $\;\;\;\;\;\frac{1}{N}\sum_{i=1}^{N}\big{\|}{\boldsymbol{\varphi}}(\boldsymbol{y}_{i,t})-\boldsymbol{x}_{i}\big{\|}^{2}\,,$ (6)
subj. to $\;\;\;\;\;{\boldsymbol{\varphi}}\in\mathscrsfs{N}_{A}\,.$
where $\boldsymbol{y}_{i,t}=t\boldsymbol{x}_{i}+\sqrt{t}\boldsymbol{g}_{i}$ for i.i.d. standard Gaussian vectors $\boldsymbol{g}_{i}$.
In applications, $\mathscrsfs{N}_{A}$ is the set of neural networks with a certain architecture $A$ and (potentially) satifying suitable bounds on the norm of the weights. We will refer to the objective in Eqs. (5), (6) as to the score matching objective or the mean square error (MSE) objective. We emphasize that, for ${\boldsymbol{\varphi}}\in\mathscrsfs{N}_{A}$ the evaluation $\boldsymbol{y}\mapsto{\boldsymbol{\varphi}}(\boldsymbol{y})$ (a.k.a. ‘inference’ in deep learning jargon) can be implemented in polynomial time. This is indeed required in order to be able to implement the generative process (4).
1.2 Summary of results
As anticipated, we construct distributions in which sampling is easy but sampling via diffusions is hard, subject to conjectured hardness of certain statistical estimation problems. Indeed the construction is fairly simple. Let $\Omega_{n,k}:=\{\boldsymbol{u}\in\{0,\pm 1/\sqrt{k}\}^{n}|\|\boldsymbol{u}\|_{0}=k\}$ be the set of $0/\pm(1/\sqrt{k})$ unit vectors with $k$ nonzero entries (the notation $\|\boldsymbol{u}\|_{0}$ denotes the number of nonzeros in $\boldsymbol{u}$). We define the target distribution $\mu=\mu_{n,k}$ to be the distribution of
$\boldsymbol{x}=\boldsymbol{u}\boldsymbol{u}^{\sf T},\;\;\;\;\;\;\;\boldsymbol{u}\sim\operatorname{Unif}(\Omega_{n,k}).$ (7)
Note that $\boldsymbol{x}\in{\mathbb{R}}^{n\times n}$ is a matrix that we identify with a vector in ${\mathbb{R}}^{d}$ for $d=n^{2}$.
Sampling from the target distribution $\mu$ is trivial. Just sample a vector with entries in $\{0,1/\sqrt{k},-1/\sqrt{k}\}$ and exactly $k$ non-zero entries, and let $\boldsymbol{x}=\boldsymbol{u}\boldsymbol{u}^{\sf T}$ as above. However, diffusion sampling requires to approximate the denoiser
$\boldsymbol{m}(\boldsymbol{y},t):=\mathbb{E}\big{\{}\boldsymbol{u}\boldsymbol{u}^{\sf T}\Big{|}\boldsymbol{y}_{t}=\boldsymbol{y}\big{\}},\;\;\;\;\boldsymbol{y}_{t}=t\boldsymbol{u}\boldsymbol{u}^{\sf T}+\sqrt{t}\boldsymbol{g},$ (8)
where $\boldsymbol{g}\sim{\mathsf{N}}({\boldsymbol{0}},{\boldsymbol{I}}_{n\times n})$. In other words, it requires to estimate a sparse rank-one matrix from a single observation in Gaussian noise. This is a well studied statistical estimation problem, closely related to the hidden clique problem, sparse PCA, and the Gaussian submatrix problem. Substantial rigorous evidence supports the claim that —for certain scalings of $k$, $t$ with $n$— polynomial-time algorithms cannot approach the Bayes-optimal error (for reference, see [BIS15, MW15, CLR17, BBH18, SW22].)
Our results suggest that sampling from $\mu_{n,k}$ using diffusions fails for $k\ll n$. More precisely, we will prove that there exists $\hat{\boldsymbol{m}}:{\mathbb{R}}^{n\times n}\times{\mathbb{R}}\to{\mathbb{R}}^{n\times n}$ such that:
  1. M1.
    $\hat{\boldsymbol{m}}(\,\cdot\,)$ can be evaluated in polynomial time.
  2. M2.
    The estimation error achieved by $\hat{\boldsymbol{m}}$ (namely, $\mathbb{E}\{\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\}$) is close to the conjectured optimal estimation error achieved by polynomial-time algorithms. Hence $\hat{\boldsymbol{m}}(\,\cdot\,,t)$ will be a near minimizer of the score-matching objective (6).
  3. M3.
    Samples $\hat{\boldsymbol{x}}_{T}$ generated by the discretized diffusions (4) with drift approximation $\hat{\boldsymbol{m}}(\,\cdot\,,t)$ at some large time $T$ have distribution that is very far from the target $\mu$.
In addition, we will prove a general reduction from estimation to diffusion sampling, establishing that if accurate diffusion sampling is possible in polynomial time, then near Bayes optimal estimation of $\boldsymbol{x}$ from $\boldsymbol{y}_{t}=t\boldsymbol{x}+\sqrt{t}\boldsymbol{g}$ must also be possible in polynomial time for all $t$. The contrapositive of this statement implies that if an information-computation gap exists, then diffusion sampling is impossible in polynomial time.
The rest of the paper is organized as follows. In Section 2 we motivate our setting and assumptions, and discuss some limitations of our results. In Section 3 we state formally our results (for technical reason we state two separate results depending on the growth of $k$ with $n$.) Section 4 presents the general reduction from estimation to diffusion sampling. Proofs are presented in Section 6 and the appendices.
Notations.
Throughout the paper it will be understood that we are considering sequences of problems indexed by $n$, where $\boldsymbol{x}\in{\mathbb{R}}^{n\times n}$ and the sparsity index $k=k_{n}$ diverges as well. We write $f(n)\ll g(n)$ or $f(n)=o(g(n))$ if $f(n)/g(n)\to 0$ and $f(n)\lesssim g(n)$ or $f(n)=O(g(n))$ if $f(n)/g(n)\leq C$ for a constant $C$. Finally $f(n)=\Theta(g(n))$ or $f(n)\asymp g(n)$ if $1/C\leq f(n)/g(n)\leq C$.
Given two measures $\mu,\nu$ on a metric space $(M,{\sf dist})$, their Wasserstein-$1$ distance is
$W_{1}(\mu,\nu):=\inf_{\gamma\in\mathcal{C}(\mu,\nu)}\int_{M\times M}{\sf dist}(\boldsymbol{x},\boldsymbol{y})\,\gamma({\rm d}\boldsymbol{x},{\rm d}\boldsymbol{y})\,,$ (9)
where the infimum is over couplings of $\mu$, $\nu$. We will apply this definitions in two cases: $(i)$ $M={\mathbb{R}}^{d}$, with ${\sf dist}(\boldsymbol{x},\boldsymbol{y})=\|\boldsymbol{x}-\boldsymbol{y}\|_{2}$; $(i)$ $M=C([0,T],{\mathbb{R}}^{d})$, with ${\sf dist}(\boldsymbol{x},\boldsymbol{y})=\sup_{t\in[0,T]}\|\boldsymbol{x}(t)-\boldsymbol{y}(t)\|_{2}$. Given random variables $\boldsymbol{X}$, $\boldsymbol{Y}$ with laws ${\rm P}_{\boldsymbol{X}}$, ${\rm P}_{\boldsymbol{Y}}$, we overload notations $W_{1}(\boldsymbol{X},\boldsymbol{Y}):=W_{1}({\rm P}_{\boldsymbol{X}},{\rm P}_{\boldsymbol{Y}})$.
We write $\boldsymbol{W}\sim{\sf GOE}(n,v)$ if $\boldsymbol{W}=\boldsymbol{W}^{{\sf T}}$ is a random symmetric matrix with $(W_{ij})_{i\leq j\leq n}$ independent entries $W_{ii}\sim{\mathsf{N}}(0,2v)$, and $W_{ij}\sim{\mathsf{N}}(0,v)$ for $i<j$. For $v=1$, we use the shorter notation ${\sf GOE}(n)$. We say that $(\boldsymbol{W}_{t}:t\geq 0)$is a ${\sf GOE}(n)$ process if $\boldsymbol{W}_{t}\in{\mathbb{R}}^{n\times n}$ is a symmetric matrix with entries above and on the diagonal $(W_{t}(i,j):i<j\leq n;W_{t}(i,i)/\sqrt{2}:i\leq n;t\geq 0)$ forming a collection of $n(n+1)/2$ independent Brownian motions.
We use $C,C_{i},c_{i},\dots$ to denote absolute constants, whose value can change from line to line. We always consider sequences of sampling problem indexed by $n$, and will write (for instance) ‘there exists an estimator $\hat{\boldsymbol{m}}(\,\cdot\,)$’ instead of the more correct ‘there exists a sequence of estimators $\hat{\boldsymbol{m}}_{n}(\,\cdot\,)$ indexed by $n\in{\mathbb{N}}$.’
2 Discussion
Setting.
Our results indicate that a standard application of denoising diffusions methodology will fail to sample from $\mu_{n,k}$ despite the sampling problem in itself is easy. On the other hand, the latent structure of the distribution $\mu_{n,k}$ can be exploited to construct a better algorithm. Namely, one can use diffusions to sample $\boldsymbol{u}\sim\operatorname{Unif}(\Omega_{n,k})$ (this is easy since the posterior expectation $\boldsymbol{m}(\boldsymbol{y},t)=\mathbb{E}[\boldsymbol{u}|t\boldsymbol{u}+\sqrt{t}\boldsymbol{g}=\boldsymbol{y}]$ is easy to compute) and then generate $\boldsymbol{x}=\boldsymbol{u}\boldsymbol{u}^{{\sf T}}$. However, identifying such latent structures from data is —in general— a hard problem, both from a statistical and from a computational viewpoint. As an example, one could consider a modification of the present problem whereby $\boldsymbol{u}$ is sparse in an unknown basis.
Limitations.
We prove that there exists drifts $\hat{\boldsymbol{m}}(\,\cdot\,,t)$ that lead to poor generative sampling, despite being nearly optimal (among poly-time algorithms) in terms of the score matching objective (5). In particular, these bad drifts will be near optimal solutions of the empirical risk minimization problem (6), as long as $\mathscrsfs{N}_{A}$ is rich enough to approximate them. This happens because the class $\mathscrsfs{N}_{A}$ is a subset of the class of functions that can be evaluated in polynomial time.
We do not formally exclude the existence of drifts $\hat{\boldsymbol{m}}(\,\cdot\,,t)$ that also satisfy conditions M1 and M2 but yield good generative sampling. However if such drift existed, our results suggest that minimizing the score matching objective is not the right approach to find them.
Correct samplers violating M2.
The subtlety in the previous point can be emphasized by the following remark. If we drop condition M2, i.e. we accept drifts that are bad from the point of view of the score matching, then it is possible to construct drifts that can be evaluated in polynomial time and yield good sampling. However, these are constructions that completely ignore the prescription (2). We provide one such construction of $\hat{\boldsymbol{m}}$ here, which guarantees that both $\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{t},t)$ and $\hat{\boldsymbol{y}}_{T}/T$ are approximate samples of $\boldsymbol{x}$ (the latter only for large $T$). Furthermore, this drift $\hat{\boldsymbol{m}}$ may depend on the noise stream $(\hat{\boldsymbol{z}}_{t})$ per Eq. (4), but relies on no additional source of randomness (the proof is presented in Appendix B.)
Proposition 2.1.
Suppose that a discretized SDE per Equation (4) is generated, with step size $\Delta>0$ and noise stream $\hat{\boldsymbol{z}}_{t}\overset{i.i.d.}{\sim}{\mathsf{N}}(0,\boldsymbol{I}_{n\times n})$. Then for every $n,k$, there exists a function $\hat{\boldsymbol{m}}(\boldsymbol{y},t)=\hat{\boldsymbol{m}}(\boldsymbol{y},t;\hat{\boldsymbol{z}}_{1})$ parametrized by $\hat{\boldsymbol{z}}_{1}$ (with no additional randomness) such that: $(i)$ $\mathbb{E}[\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}]=2(1-o(1))$ uniformly for every $t\geq 0$ (sub-optimal score-matching error); $(ii)$ $W_{1}(\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{\ell\Delta},\ell\Delta),\boldsymbol{x})=0$ for all $\ell\geq 0$ ($\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{t},t)$ is an approximate sample of $\boldsymbol{x}$ at every point); $(iii)$ $\lim_{\ell\to\infty}W_{1}(\hat{\boldsymbol{y}}_{\ell\Delta}/(\ell\Delta),\boldsymbol{x})=0$ ($\hat{\boldsymbol{y}}_{t}/t$ is an approximate sample of $\boldsymbol{x}$ at large time).
Again, it is crucial to observe that the drift constructed in this proposition is artificial and has very poor value of the score-matching objective.
Hardness assumption.
Condition M2 refers to the conjectured optimal estimation error achieved by polynomial-time algorithms because lower bounds on this error do necessarily rely either on complexity-theoretic assumptions or on optimality or certain restricted classes of algorithms.
Related work.
A number of groups proved positive results on diffusion sampling. Among others, [AMS22, CCL+23, MW23, LLT23, BDBDD23] provide reductions from diffusion sampling to score estimation, while [CLL23, SCK23, MW25, LHW24] establish end-to-end guarantees for classes of distributions $\mu$.
The computational bottleneck that we study here has been observed before in the context of certain Gibbs measures and Bayes posterior distributions [GDKZ24, AMS23, HMP24], and random constraint satisfaction problems [MRTS07, RTS09] (the latter papers use however sequential sampling rather than diffusion sampling).
Our work provides a simpler example and a rigorous formalization of the latter line of work. (Notice that despite superficial similarities, our construction is different from the examples of [GDKZ24, AMS23, HMP24], and allows us to build existing hardness results.)
3 Main results
For technical reasons, we will state two separate results for the probability distribution $\mu=\mu_{n,k}$ described in the introduction, depending on the scaling of $k$ with $n$: in Section 3.1 we assume $\sqrt{n}\ll k\ll n$; while in Section 3.3 we assume $k\ll\sqrt{n}$. The reason for this distinction is that the nature of the estimator $\hat{\boldsymbol{m}}$ changes at the threshold $k\asymp\sqrt{n}$.
In what follows $(\boldsymbol{x},\boldsymbol{y}_{t})$ will always be distributed according to the ideal diffusion process of Eq. (3), which also satisfies Eq. (2), and $\hat{\boldsymbol{y}}_{t}$ will denote the process generated with the implemented procedure (4). A crucial role will be played by the following algorithmic threshold
$t_{\mbox{\tiny\rm alg}}(n,k):=\begin{cases}k^{2}\log\left(\dfrac{n}{k^{2}}\right)&\text{ if $k\ll\sqrt{n}$}\\ \dfrac{n}{2}&\text{ if $\sqrt{n}\ll k\ll n$}\end{cases}$ (10)
As discussed below, see Conjecture 3.2, it is expected that no polynomial-time algorithm can estimate $\boldsymbol{x}$ significantly better than the null estimator $\hat{\boldsymbol{m}}_{\mbox{\tiny\rm null}}=\boldsymbol{0}$ for ${\underline{k}}_{n}\leq k\ll n$, with ${\underline{k}}_{n}\ll n$ a suitable sequence.
We also note that since $\|\boldsymbol{x}\|_{F}=1$ almost surely under the target distribution $\boldsymbol{x}\sim\mu$, the Bayes optimal denoiser $\boldsymbol{m}(\,\cdot\,)$ does not depend explicitly on $t$ and we will therefore use the notation $\boldsymbol{m}(\boldsymbol{y}_{t})$.
3.1 Moderately sparse regime: $\sqrt{n}\ll k\ll n$
We next state our main result in the moderately sparse regime $\sqrt{n}\ll k\ll n$. Given an arbitrary polynomial time computable drift $\hat{\boldsymbol{m}}_{0}$, we construct a different poly-time drift $\hat{\boldsymbol{m}}$, with nearly equal score matching objective and yet incorrect sampling. In the next subsection we state a related result that establishes incorrect sampling from an explicit denoiser with near-optimal denoising properties.
Assumption 1 (Small norm below threshold).
Let $\boldsymbol{y}_{t}=t\boldsymbol{x}+\boldsymbol{B}_{t}$, and assume $\sqrt{n}\ll k\ll n$. Then, there exists (known) constants $\gamma,\varepsilon\in(0,1)$, independent of $n$ such that, for every $D>0$,
$\int_{0}^{(1-\gamma)t_{\mbox{\tiny\rm alg}}}\mathbb{P}\big{(}\|\hat{\boldsymbol{m}}_{0}(\boldsymbol{y}_{t},t)\|\geq 1-\varepsilon\big{)}\,{\rm d}t=O(n^{-D})\,.$
We expect this condition to be satisfied by any reasonable polynomial time algorithm, under the well-accepted Conjecture 3.2 below on information-computation gaps. Indeed, the conjecture states that $\inf_{0\leq t\leq(1-\gamma)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\left[\|\hat{\boldsymbol{m}}_{0}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\right]\geq 1-o_{n}(1)$. A simple calculation shows that any polytime estimator matching this error must also satisfy $\mathbb{E}\{\|\hat{\boldsymbol{m}}_{0}(\boldsymbol{y}_{t},t)\|^{2}\}=o_{n}(1)$ (because otherwise $c_{t}\hat{\boldsymbol{m}}_{0}(\,\cdot\,)$ would violate the lower bound for a suitable deterministic $c_{t}$).
Theorem 1.
Assume $\sqrt{n}\ll k\ll n$, and note that in this case $t_{\mbox{\tiny\rm alg}}(n,k):=n/2$ per Eq. (10). Let $\hat{\boldsymbol{m}}_{0}:{\mathbb{R}}^{n\times n}\times{\mathbb{R}}\to{\mathbb{R}}^{n\times n}$ be an arbitrary poly-time algorithm such that $\sup_{\boldsymbol{y},t}\|\hat{\boldsymbol{m}}_{0}(\boldsymbol{y},t)\|_{F}\leq 1$ and Assumption 1 holds. Then there exists an estimator $\hat{\boldsymbol{m}}:{\mathbb{R}}^{n\times n}\times{\mathbb{R}}\to{\mathbb{R}}^{n\times n}$ such that:
  • M1.
    $\hat{\boldsymbol{m}}(\,\cdot\,)$ can be evaluated in polynomial time.
  • M2.
    If $\boldsymbol{y}_{t}=t\boldsymbol{x}+\boldsymbol{B}_{t}$ is the true diffusion (equivalently given by Eq. (1)), then, for every $D>0$,
    $\int_{0}^{\infty}\mathbb{E}[\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\hat{\boldsymbol{m}}_{0}(\boldsymbol{y}_{t},t)\|^{2}]\,{\rm d}t=O(n^{-D})\,.$
  • M3.
    There exists $\delta=o_{n}(1)$ such that, for any step size $\Delta=\Delta_{n}>0$, we have
    $\inf_{t\in\mathbb{N}\cdot\Delta,t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}}W_{1}(\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{t},t),\boldsymbol{x})\geq 1-o_{n}(1)\,.$ (11)
In order to connect this theorem with the general discussion in the introduction, we recall two facts from the literature on submatrix estimation: $(i)$ The ideal posterior mean estimator $\boldsymbol{m}(\boldsymbol{y}_{t})$ achieves small mean square error in a large interval below $t_{\mbox{\tiny\rm alg}}$ (Proposition 3.1); $(ii)$ No polynomial-time estimator is expected to perform substantially better than the null estimator below $t_{\mbox{\tiny\rm alg}}$ (Conjecture 3.2).
To provide more detail on point $(i)$, we state below a characterization of the Bayes optimal error. The proof is essentially the same (indeed simpler) as the main result in [BIS15], which considers the case of asymmetric matrices. An alternative proof (for $k\leq n^{a}$, $a<5/6$) also follows from [BMR20].
Proposition 3.1 (Modification of [BIS15]).
Let $\boldsymbol{m}(\,\cdot\,)$ be the posterior mean estimator in the Gaussian submatrix problem (8). Assume $1\ll k\ll n$, and define $t_{\mbox{\tiny\rm Bayes}}(n,k):=2k\log(n/k)$. Then, for any $\delta>0$, we have
$\inf_{t\leq(1-\delta)t_{\mbox{\tiny\rm Bayes}}}\mathbb{E}\big{\{}\|\boldsymbol{m}(\boldsymbol{y}_{t})-\boldsymbol{x}\|^{2}\}=1-o_{n}(1)\,,\;\;\;\;\;\;\sup_{t\geq(1+\delta)t_{\mbox{\tiny\rm Bayes}}}\mathbb{E}\big{\{}\|\boldsymbol{m}(\boldsymbol{y}_{t})-\boldsymbol{x}\|^{2}\}=o_{n}(1)\,.$
Several groups have proved hardness results for estimating $\boldsymbol{u}$ in the model considered here, for $t$ sufficiently smaller than the spectral threshold $n/2$. In particular [BBH18] establishes hardness conditional on hardness on the planted clique problem for $t\leq n^{1-\varepsilon}$, $\varepsilon>0$, [HKP+17] proved failure of sum-of-squares relaxations also for $t\leq n^{1-\varepsilon}$, [SW22] proved hardness for low-degree polynomials for the related hidden submatrix problem for $t\leq cn$, $c>0$ a sufficiently small constant. Finally, [KWB19] prove hardness for low-degree polynomials algorithms when $t\leq(1-\varepsilon)n/2$, albeit their results require $k$ to be of order $n$.
Conjecture 3.2.
There exists $\underline{k}_{n}\ll n$ such that the following holds for any $k=k_{n}$, with $\underline{k}_{n}\leq k_{n}\ll n$. Let $\{\hat{\boldsymbol{m}}_{n}\}_{n\geq 1}$, $\hat{\boldsymbol{m}}_{n}:{\mathbb{R}}^{n\times n}\times{\mathbb{R}}\to{\mathbb{R}}^{n\times n}$ be any sequence of polynomial time algorithms indexed by $n$ (namely $\hat{\boldsymbol{m}}_{n}$ can be evaluated in time bounded by $n^{C}$ for some constant $C$). Then for any $\delta>0$, we have
$\inf_{t\leq(1-\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}_{n}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\}$ $\geq 1-o_{n}(1)\,.$ (12)
Summarizing, in the moderately sparse regime, for $2k\log(n/k)\ll t\ll n$ the Bayes optimal estimator can estimate the signal $\boldsymbol{x}$ accurately, but we expect that no polynomial-time algorithm can achieve the same.
Remark 3.1.
It makes sense to assume that $\|\hat{\boldsymbol{m}}_{0}(\cdot,\cdot)\|_{F}\leq 1$. Since $\boldsymbol{x}\sim\mu_{n,k}$ and $\text{supp}(\mu_{n,k})\subset\{\boldsymbol{a}:\|\boldsymbol{a}\|_{F}\leq 1\}$ and the latter is a convex set, projecting any estimator $\hat{\boldsymbol{m}}_{0}$ onto this set yields a smaller meansquare error.
3.2 Moderately sparse regime: A concrete example
Algorithm 1 Submatrix Estimation Algorithm (moderately sparse regime)
1:  Input: Data $\boldsymbol{y}_{t}$; time $t$; parameter $\varepsilon$
2:  Output: Estimate of $\boldsymbol{x}$: $\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)$
3:  If $t\geq t_{\mbox{\tiny\rm alg}}$, continue; otherwise return $\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)={\boldsymbol{0}}$
4:  Symmetrize: $\boldsymbol{A}_{t}=(\boldsymbol{y}_{t}+\boldsymbol{y}_{t}^{\sf T})/(2\sqrt{t})$
5:  If $t\geq n^{2}$ and $\lambda_{1}(\boldsymbol{A}_{t})\leq\sqrt{t}/2$, return $\boldsymbol{0}$; otherwise continue
6:  Compute top eigenvector of $\boldsymbol{A}_{t}$, denoted if by $\boldsymbol{v}_{t}$
7:  Compute $\hat{S}$ by
$\hat{S}:=\Big{\{}i\in[n]:\;|v_{t,i}|\geq\frac{\varepsilon}{\sqrt{k}}\Big{\}}\,;$ (13)
8:  Compute vector $\boldsymbol{w}$ such that $w_{i}=\operatorname{sign}(v_{t,i})\boldsymbol{1}_{i\in\hat{S}}$
9:  return  $\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t):=\boldsymbol{w}\boldsymbol{w}^{{\sf T}}/|\hat{S}|$ if $|\hat{S}|\geq k/2$; otherwise return $\boldsymbol{0}$
Since Theorem 1 is somewhat abstract, we complement it with an explicit example. Namely, we prove that a modification of a standard spectral estimator, while achieving near optimal estimation error (among polytime algorithms) fails to generate samples from the correct distribition.
Theorem 2.
Assume $\sqrt{n}\ll k\ll n$, and note that in this case $t_{\mbox{\tiny\rm alg}}(n,k):=n/2$ per Eq. (10). Then the estimator $\hat{\boldsymbol{m}}:{\mathbb{R}}^{n\times n}\times{\mathbb{R}}\to{\mathbb{R}}^{n\times n}$ defined in Algorithm 1 satisfies the following:
  • M1.
    $\hat{\boldsymbol{m}}(\,\cdot\,)$ can be evaluated in polynomial time.
  • M2.
    For any $\delta>0$, there exists $c=c(\delta)$, $C=C(\delta)$ such that
    $\inf_{t\leq(1-\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\big{\}}=1-o_{n}(1)\,,\;\;\;\;\;\;\sup_{t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\big{\}}\leq C\,e^{-cn/k}\,.$
  • M3.
    For any $\Delta>0$,
    $\inf_{t\in{\mathbb{N}}\cdot\Delta}W_{1}(\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{t},t),\boldsymbol{x})=1-o_{n}(1)\,.$
As mentioned, Algorithm 1 uses a spectral approach. We compute the leading eigenvector of (the symmetrized version of) $\boldsymbol{y}_{t}$, call it $\boldsymbol{v}_{t}\in{\mathbb{R}}^{n}$. We then estimate the support $S$ of the latent rank-one matrix using the largest entries of $\boldsymbol{v}_{t}$.
3.3 Very sparse regime: $k\ll\sqrt{n}$
Algorithm 2 Submatrix Estimation Algorithm (very sparse regime)
1:  Input: Data $\boldsymbol{y}_{t}$; time $t$; parameters $s,\varepsilon$
2:  Output: Estimate of $\boldsymbol{x}$: $\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)$
3:  Let $\boldsymbol{g}_{t}\sim{\mathsf{N}}(0,t{\boldsymbol{I}}_{n\times n})$ and compute
$\boldsymbol{y}_{t,+}:=\boldsymbol{y}_{t}+\sqrt{\varepsilon}\boldsymbol{g}_{t},,\;\;\;\;\boldsymbol{y}_{t,-}:=\boldsymbol{y}_{t}-\sqrt{\frac{1}{\varepsilon}}\boldsymbol{g}_{t}\,.$ (14)
4:  Symmetrize: $\boldsymbol{A}_{t,+}=(\boldsymbol{y}_{t,+}+\boldsymbol{y}_{t,+}^{\sf T})/(2\sqrt{t})$, $\boldsymbol{A}_{t,-}=(\boldsymbol{y}_{t,-}+\boldsymbol{y}_{t,-}^{\sf T})/(2\sqrt{t})$
5:  Compute top eigenvector of $\eta_{s}(\boldsymbol{A}_{t,+})$, denoted if by $\boldsymbol{v}_{t}$
6:  If $t\geq t_{\mbox{\tiny\rm alg}}\vee 1$ and $\lambda_{1}\left(\eta_{s}\left(\boldsymbol{A}_{t,+}\right)\right)>k+\dfrac{\sqrt{t}}{s}$, continue; otherwise return $\hat{\boldsymbol{m}}(\boldsymbol{y},t):=\boldsymbol{0}$
7:  Compute the vector $\hat{\boldsymbol{v}}_{t}:=\boldsymbol{A}_{t,-}\boldsymbol{v}_{t}$
8:  Let $\hat{S}$ be the set of $k$ indices $i$ of largest values of $|\hat{v}_{t,i}|$, and compute vector $\boldsymbol{w}$ such that $w_{i}=\operatorname{sign}(\hat{v}_{t,i})\boldsymbol{1}_{i\in\hat{S}}$
9:  return  $\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t):=\boldsymbol{1}_{\hat{S}}\boldsymbol{1}_{\hat{S}}^{{\sf T}}/k$
Theorem 3.
Assume $(\log n)^{5/2}\lesssim k\ll\sqrt{n}$, and note that in this case $t_{\mbox{\tiny\rm alg}}(n,k)=k^{2}\log(n/k^{2})$, per Eq. (10). Then the randomized estimeator $\hat{\boldsymbol{m}}:{\mathbb{R}}^{n\times n}\times{\mathbb{R}}\to{\mathbb{R}}^{n\times n}$ of Algorithm 2 satisfies the following:
  • M1.
    $\hat{\boldsymbol{m}}(\,\cdot\,)$ can be evaluated in polynomial time.
  • M2.
    For any $\delta>0$, $D>0$, there exists $C(\delta,D)>0$ such that:
    $\inf_{t\leq(1-\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\big{\}}=1-o_{n}(1)\,,\;\;\;\;\;\;\sup_{t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\big{\}}\leq C(\delta,D)n^{-D}\,.$
  • M3.
    For any $\Delta>0$,
    $\inf_{t\in{\mathbb{N}}\cdot\Delta}W_{1}(\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{t},t),\boldsymbol{x})=1-o_{n}(1)\,.$
The pseudocode for the estimator $\hat{\boldsymbol{m}}(\,\cdot\,)$ that is constructed in the above is given as Algorithm 2. This is based on a standard approach in the literature [DM16, CLR17], with some modifications to allow for its analysis in the diffusion setting. The main steps are as follows:
  • Perform Gaussian data splitting of $\boldsymbol{y}_{t}$ into $\boldsymbol{y}_{t,+}$, $\boldsymbol{y}_{t,-}$, see Eq. (14), with most of the information preserved in $\boldsymbol{y}_{t,+}$.
  • Use entrywise soft thresholding $\eta_{s}(x)=(|x|-s)_{+}\operatorname{sign}(x)$ to reduce the noise in the symmetrized version of $\boldsymbol{y}_{t,+}$.
  • Compute a first estimate of the latent vector $\boldsymbol{1}_{S}$ by the principal eigenvector of the above matrix.
  • Refine this estimate using the remaining information $\boldsymbol{y}_{t,-}$.
We point out that Proposition 3.1 remains true in the regime $\sqrt{n}\ll k\ll n$, and hence we observe a gap between $t_{\mbox{\tiny\rm alg}}(n,k)$ and $t_{\mbox{\tiny\rm Bayes}}(n,k)$ in this regime as well. However, we are not aware of matching hardness results in this regime.
4 Reduction of estimation to diffusion-based sampling
In order to further support and clarify our main conclusions, we present two simple results showing that estimation in the Gaussian noise model is reducible to diffusion sampling. Informally, if diffusion sampling can be performed in polynomial time with sufficient accuracy, then we can sample in polynomial time from the posterior distribution of $\boldsymbol{x}\sim\mu$ given noisy information $\boldsymbol{y}=\boldsymbol{x}+\sigma\boldsymbol{g}$ (where $\boldsymbol{g}$ is standard Gaussian noise independent of $\boldsymbol{x}$).
Here we consider a general sampling problem for $\boldsymbol{x}\sim\mu$ a probability distribution in ${\mathbb{R}}^{d}$. For the sake of simplicity, we will assume $\mu$ to be supported on the unit sphere ${\mathbb{S}}^{d-1}=\{\boldsymbol{x}:\|\boldsymbol{x}\|=1\}$. Our assumptions will be about the accuracy of the denoising diffusion process. We denote by ${\rm P}_{\boldsymbol{y}}^{T}$ the law of $(\boldsymbol{y}_{t})_{0\leq t\leq T}$ where $\boldsymbol{y}_{t}$ given by Eq. (1) and by ${\rm P}_{\hat{\boldsymbol{y}}}^{T,\Delta}$ the law of $(\hat{\boldsymbol{y}}_{t})_{0\leq t\leq T}$, which is the discretized diffusion trajectory defined in (4) (interpolated linearly outside ${\mathbb{N}}\cdot\Delta$).
It is further useful to define $\overline{\rm P}_{\hat{\boldsymbol{y}}}^{T,\Delta}$ to be the law of the the solution following SDE
${\rm d}\hat{\boldsymbol{y}}_{t}=\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{\lfloor t\rfloor_{\Delta}},\lfloor t\rfloor_{\Delta})\,{\rm d}t+{\rm d}\boldsymbol{B}_{t}\,,$ (15)
where $\lfloor t\rfloor_{\Delta}:=\max\{s\in{\mathbb{N}}\cdot\Delta:s\leq t\}$. It is easy to see that this definition interpolates $(\hat{\boldsymbol{y}}_{t}:t\in{\mathbb{N}}\cdot\Delta)$ defined in (4)
We state first an easier form of the reduction.
Theorem 4.
Assume that $\hat{\boldsymbol{m}}(\,\cdot\,,\,\cdot\,)$ has complexity $\chi$ and that for any $T\leq\theta d$, $D_{\mbox{\tiny\rm KL}}(\overline{\rm P}_{\hat{\boldsymbol{y}}}^{T,\Delta}\|{\rm P}_{\boldsymbol{y}}^{T})\leq\varepsilon$ (where $\overline{\rm P}_{\hat{\boldsymbol{y}}}^{T,\Delta}$ is the continuous time process obtained by Brownian-linear interpolation of Eq. (4)).
Then for any $\sigma>0$ there exists an algorithm with complexity $O(\chi\cdot T/\Delta)$, that takes as input $\boldsymbol{y}=\boldsymbol{x}+\sigma\boldsymbol{g}$, $(\boldsymbol{x},\boldsymbol{g})\sim\mu\otimes{\mathsf{N}}(0,{\boldsymbol{I}})$, and outputs $\hat{\boldsymbol{x}}$, such that
$\mathbb{E}\|{\rm P}_{\boldsymbol{x}|\boldsymbol{y}}-{\rm P}_{\hat{\boldsymbol{x}}|\boldsymbol{y}}\|_{\mbox{\tiny\rm TV}}\leq\sqrt{2\varepsilon}+\varepsilon_{0}(\theta)=:\overline{\varepsilon}\,,$ (16)
where $\varepsilon_{0}(\theta):=\mathbb{E}\|{\rm P}_{\boldsymbol{x}|\boldsymbol{y}}-{\mathsf{N}}(\boldsymbol{0},(\theta d)^{-1}\boldsymbol{I}_{d})*{\rm P}_{\boldsymbol{x}|\boldsymbol{y}}\|_{\mbox{\tiny\rm TV}}$ is the expected TV distance between ${\rm P}_{\boldsymbol{x}|\boldsymbol{y}}$ and the convolution of ${\rm P}_{\boldsymbol{x}|\boldsymbol{y}}$ with a Gaussian with variance $1/(\theta d)$. As a consequence, there exists a randomized algorithm $\hat{\boldsymbol{m}}_{+}$ with complexity $(N\chi\cdot T/\Delta)$ that approximates the posterior expectation, namely
$\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}_{+}(\boldsymbol{y})-\mathbb{E}(\boldsymbol{x}|\boldsymbol{y})\|^{2}\big{\}}\leq 2\overline{\varepsilon}+\frac{2}{N}\,.$ (17)
Proof.
The algorithm consists in running the discretized diffusion (4) with initialization $\hat{\boldsymbol{y}}_{t_{0}}=\boldsymbol{y}/\sigma^{2}$ at $t=t_{0}:=1/\sigma^{2}$. To avoid notational burden, we will assume $(T-t_{0})/\Delta$ to be an integer. Let $\hat{\boldsymbol{y}}_{t_{0}}^{*}$ be generated by the discretized diffusion with initialization at $\hat{\boldsymbol{y}}_{0}$ at $t=0$. Note that the distribution of $\hat{\boldsymbol{y}}_{t_{0}}$ is the same as the one of $t_{0}\boldsymbol{x}+\sqrt{t}\boldsymbol{g}$ and hence by Assumption $(b)$, and Pinsker’s inequality
$\|{\rm P}_{\hat{\boldsymbol{y}}_{t_{0}}}-{\rm P}_{\hat{\boldsymbol{y}}^{*}_{t_{0}}}\|_{\mbox{\tiny\rm TV}}\leq\sqrt{\frac{1}{2}D_{\mbox{\tiny\rm KL}}({\rm P}_{\hat{\boldsymbol{y}}^{*}_{t_{0}}}\|{\rm P}_{\hat{\boldsymbol{y}}_{t_{0}}})}\leq\sqrt{\frac{1}{2}D_{\mbox{\tiny\rm KL}}(\overline{\rm P}_{\hat{\boldsymbol{y}}}^{T,\Delta}\|{\rm P}_{\boldsymbol{y}}^{T})}\leq\sqrt{\frac{\varepsilon}{2}}\,.$ (18)
Hence $\hat{\boldsymbol{y}}_{t_{0}}$, $\hat{\boldsymbol{y}}^{*}_{t_{0}}$ can be coupled so that $\mathbb{P}(\hat{\boldsymbol{y}}_{t_{0}}\neq\hat{\boldsymbol{y}}^{*}_{t_{0}})\leq\sqrt{\varepsilon/2}$.
We extend this to a coupling of $(\hat{\boldsymbol{y}}^{*}_{t})_{t_{0}\leq t\leq T}$ and $(\hat{\boldsymbol{y}}_{t})_{t_{0}\leq t\leq T}$ in the obvious way: we generate the two trajectories according to the discretized diffusion (4) with the same randomness $\hat{\boldsymbol{z}}_{t}$. Therefore $\mathbb{P}(\hat{\boldsymbol{y}}_{T}\neq\hat{\boldsymbol{y}}^{*}_{T})\leq\sqrt{\varepsilon/2}$. Another application of the assumption $D_{\mbox{\tiny\rm KL}}(\overline{\rm P}_{\hat{\boldsymbol{y}}}^{T,\Delta}\|{\rm P}_{\boldsymbol{y}}^{T})\leq\varepsilon$ and Pinsker’s inequality yields $\mathbb{P}(\boldsymbol{y}_{T}\neq\hat{\boldsymbol{y}}^{*}_{T})\leq\sqrt{\varepsilon/2}$, for $\boldsymbol{y}_{T}\stackrel{{\scriptstyle{\rm d}}}{{=}}T\boldsymbol{x}+\sqrt{T}\boldsymbol{g}^{\prime}$ with $(\boldsymbol{x},\boldsymbol{g}^{\prime})\sim\mu\otimes{\mathsf{N}}(\boldsymbol{0},\boldsymbol{I})$. We conclude by triangle inequality $\mathbb{P}(\boldsymbol{y}_{T}\neq\hat{\boldsymbol{y}}_{T})\leq 2\sqrt{\varepsilon/2}$, which coincides with the claim (16).
Finally, Eq. (17) follows by generating $N$ i.i.d. copies $\hat{\boldsymbol{x}}_{1},\dots,\hat{\boldsymbol{x}}_{N}$ using the above procedure, and letting $\hat{\boldsymbol{m}}(\boldsymbol{y})$ be their empirical average. ∎
The next statement makes a weaker assumption on the accuracy of the diffusion sampler (transportation instead of KL distance), but in exclnage assumes the approximate drift $\hat{\boldsymbol{m}}$ to be Lipschitz. We note that ${\rm Lip}(\boldsymbol{m}(\,\cdot\,,t))=\sup_{\boldsymbol{y}}\|{\rm Cov}(\boldsymbol{x}|\boldsymbol{y}_{t}=\boldsymbol{y})\|_{{\mbox{\rm\tiny op}}}$, and the latter is of $O(1/d)$ (for instance) if the coordinates of $\boldsymbol{x}$ are weakly dependent under the posterior.
Theorem 5.
Assume that $\hat{\boldsymbol{m}}(\,\cdot\,,\,\cdot\,)$ has computational complexity $\chi$ and satisfies the following:
  1. $(a)$
    For every $t\geq 1/\sigma^{2}$, $\boldsymbol{y}\mapsto\hat{\boldsymbol{m}}(\boldsymbol{y},t)$ is $L/d$-Lipschitz.
  2. $(b)$
    There is a stepsize $\Delta$ such that $W_{1}({\rm P}_{\hat{\boldsymbol{y}}}^{T,\Delta},{\rm P}_{\boldsymbol{y}}^{T})\leq\varepsilon$ for any $T\leq\theta d$.
Then for any $\sigma>0$ there exists an algorithm with complexity $O(\chi\cdot T/\Delta)$, that takes as input $\boldsymbol{y}=\boldsymbol{x}+\sigma\boldsymbol{g}$, $(\boldsymbol{x},\boldsymbol{g})\sim\mu\otimes{\mathsf{N}}(0,{\boldsymbol{I}})$, and outputs $\hat{\boldsymbol{x}}$, such that
$\mathbb{E}_{\boldsymbol{y}}W_{1}({\rm P}_{\boldsymbol{x}|\boldsymbol{y}},{\rm P}_{\hat{\boldsymbol{x}}|\boldsymbol{y}})\leq 2e^{\theta L}\varepsilon+\frac{1}{\sqrt{\theta}}=:\overline{\varepsilon}\,.$ (19)
As a consequence, Eq. (17) holds also in this case with the new definition of $\overline{\varepsilon}$.
5 Proof of Theorem 1
5.1 Auxiliary lemmas
We will use the following lemmas, whose proofs are deferred to Appendix E.
Lemma 5.1.
Let $\boldsymbol{W}\sim{\sf GOE}(n,1/2)$, and $C>\sqrt{2}$ some positive constant. Then we have
$\mathbb{P}\left(\max_{\boldsymbol{v}\in\Omega_{n,k}}|\langle\boldsymbol{v},\boldsymbol{W}\boldsymbol{v}\rangle|\geq C\sqrt{\log{\binom{n}{k}}}\right)\leq 2\binom{n}{k}^{-C^{2}/2+2}\,.$
We state the following non-asymptotic result from [Pen12].
Lemma 5.2 (Theorem 3.1, [Pen12].).
Let $\boldsymbol{y}=\theta\boldsymbol{u}\boldsymbol{u}^{{\sf T}}+{\sf GOE}(n,1/n)$, and denote by $\lambda_{1}(\boldsymbol{y})$ the top eigenvalue of $\boldsymbol{y}$. Letting $\xi(\theta):=\theta+\theta^{-1}$, the following holds for every $x\geq 0$ and $\theta>1$:
$\mathbb{P}\left(\lambda_{1}(\boldsymbol{y})\leq\xi(\theta)-x-\frac{2}{n}\right)\leq\exp\left(-\dfrac{(n-1)(\theta-1)^{4}}{16\theta^{2}}\right)+8\exp\left(-\dfrac{1}{4}\cdot\dfrac{(n-1)(\theta-1)^{5}x^{2}}{(\theta+1)^{3}}\right)\,.$
In Appendix E, we will use the last lemma to prove the bound below.
Lemma 5.3 (Alignment bound).
Let $\boldsymbol{y}=\theta\boldsymbol{u}\boldsymbol{u}^{{\sf T}}+{\sf GOE}(n,1/n)$, where $\theta=\sqrt{1+\delta}$. Let $\boldsymbol{v}_{1}$ be the top eigenvector of $\boldsymbol{y}$. Then, there exist constants $C,c>0$ such that the following holds for any $\delta=\delta_{n}$ with $n^{-c}\ll\delta\ll 1$ for some $c>0$ small enough:
$\mathbb{P}\left(|\langle\boldsymbol{v}_{1},\boldsymbol{u}\rangle|\leq c\delta^{2}\right)\leq C\,e^{-cn^{1/3}}\,.$ (20)
We also use the following lemma, which is implied in the proof of Lemma 6.4.
Lemma 5.4.
Let $\boldsymbol{g}\sim{\mathsf{N}}(\boldsymbol{0},\boldsymbol{I}_{n})$ and define the set
${\mathcal{L}}(\boldsymbol{g};C)=\left\{i:1\leq i\leq n,|g_{i}|\geq C\sqrt{\log(n/k)}\right\}\,.$
Assume $\sqrt{n}\ll k\ll n$. Then, for any $C$ large enough, there exists $C_{*}$ such that
$\mathbb{P}\left(|{\mathcal{L}}(\boldsymbol{g};C)|\geq\Big{(}\sqrt{k}\vee\frac{k^{2}}{n}\Big{)}\right)\leq C_{*}e^{-n^{1/4}}\,.$
5.2 Proof of Theorem 1
Let $\delta=o_{n}(1)$ be a parameter to be chosen later and recall that $t_{\mbox{\tiny\rm alg}}=n/2$. We define $\hat{\boldsymbol{m}}$ as follows:
$\hat{\boldsymbol{m}}(\boldsymbol{y},t)=\begin{cases}\hat{\boldsymbol{m}}_{0}(\boldsymbol{y},t)\boldsymbol{1}_{\|\hat{\boldsymbol{m}}_{0}(\boldsymbol{y},t)\|\leq 1-\varepsilon}&\text{ if $t\leq(1-\gamma)t_{\mbox{\tiny\rm alg}}$,}\\ \hat{\boldsymbol{m}}_{0}(\boldsymbol{y},t)&\text{ if $(1-\gamma)t_{\mbox{\tiny\rm alg}}<t<(1+\delta)t_{\mbox{\tiny\rm alg}}$,}\\ \hat{\boldsymbol{m}}_{0}(\boldsymbol{y},t)\phi_{1}(\boldsymbol{y},t)&\text{ if$(1+\delta)t_{\mbox{\tiny\rm alg}}\leq t<n^{4}$,}\\ \hat{\boldsymbol{m}}_{0}(\boldsymbol{y},t)\phi_{2}(\boldsymbol{y},t)&\text{ if$t\geq n^{4}$,}\end{cases}$
where $\phi_{1},\phi_{2}:{\mathbb{R}}^{n\times n}\times{\mathbb{R}}_{\geq 0}\to\{0,1\}$ are defined below and $\gamma$ is given in Assumption 1. It will be clear from the constructions below that $\phi_{1},\phi_{2}$ can be evaluated in polynomial time.
In order to prove claim M2, we need to bound $J_{n}(0,\infty)$, whereby, for $t_{a}\leq t_{b}$,
$J_{n}(t_{a},t_{b}):=\int_{t_{a}}^{t_{b}}\mathbb{E}\left[\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\hat{\boldsymbol{m}}_{0}(\boldsymbol{y}_{t},t)\|^{2}\right]{\rm d}t$
Setting $t_{1}=(1+\delta)t_{\mbox{\tiny\rm alg}}$ and $t_{2}=n^{4}$, we write
$J_{n}(0,\infty)=J_{n}(0,t_{1})+J_{n}(t_{1},t_{2})+J_{n}(t_{2},\infty)\,,$ (21)
and will bound each of the three terms separately.
Bounding $J_{n}(0,t_{1})$.
By Assumption 1, we know that
$J_{n}(0,t_{1})\leq\int_{0}^{(1-\gamma)t_{0}}\mathbb{P}(\|\hat{\boldsymbol{m}}_{0}(\boldsymbol{y}_{t},t)\|>1-\varepsilon)\,{\rm d}t=O(n^{-D})\,,$ (22)
for every $D>0$.
Bounding $J_{n}(t_{1},t_{2})$.
For a matrix $\boldsymbol{y}\in{\mathbb{R}}^{n\times n}$ and time point $t$, we define $\phi_{1}(\boldsymbol{y},t)$ according to the following procedure:
  1. 1.
    Compute sthe symmetrized matrix $\boldsymbol{A}=(1/2)(\boldsymbol{y}+\boldsymbol{y}^{{\sf T}})$;
  2. 2.
    Compute its top eigenvector $\boldsymbol{v}$ (choose at random if this is not unique).
  3. 3.
[…]
  4. 4.
    Compute the test statistic $s:=\langle\hat{\boldsymbol{v}},\boldsymbol{A}\hat{\boldsymbol{v}}\rangle$, and return $1$ if $s\geq\beta t$; 00 otherwise.
Here $\beta\in(0,1)$ is a fixed constant to be chosen later.
We will show that $\phi_{1}(\boldsymbol{y}_{t},t)=1$ with overwhelming probability for the true model $\boldsymbol{y}_{t}=t\boldsymbol{x}+\boldsymbol{B}_{t}$. Define
$\boldsymbol{A}_{t}=\dfrac{\boldsymbol{y}_{t}+\boldsymbol{y}_{t}^{{\sf T}}}{2}=t\boldsymbol{u}\boldsymbol{u}^{{\sf T}}+\boldsymbol{W}_{t}$
where we recall that $\boldsymbol{u}\sim\operatorname{Unif}(\Omega_{n,k})$ and $\boldsymbol{W}_{t}$ is a ${\sf GOE}(n)$ process. Let $\boldsymbol{v}_{t},\hat{\boldsymbol{v}}_{t}$ be the top eigenvector and thresholded vector of $\boldsymbol{A}_{t}$, respectively.
We have
$s_{t}:=\langle\hat{\boldsymbol{v}}_{t},\boldsymbol{A}_{t}\hat{\boldsymbol{v}}_{t}\rangle=t\cdot\langle\boldsymbol{u},\hat{\boldsymbol{v}}_{t}\rangle^{2}+\langle\hat{\boldsymbol{v}}_{t},\boldsymbol{W}_{t}\hat{\boldsymbol{v}}_{t}\rangle\,.$ (23)
Using Lemma 5.1, we know that $|\langle\hat{\boldsymbol{v}}_{t},\boldsymbol{W}\hat{\boldsymbol{v}}_{t}\rangle|\leq 4\sqrt{k\log(n/k)}\cdot\sqrt{t}$ with probability at least $1-\binom{n}{k}^{-6}$, say. Further
$\boldsymbol{v}_{t}=\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle\boldsymbol{u}+\sqrt{1-\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle^{2}}\,\boldsymbol{w}\,,$
where $\boldsymbol{w}$ is a uniformly random unit vector orthogonal to $\boldsymbol{u}$, independent of $\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle$. Alternatively, there exists $\boldsymbol{g}\sim{\mathsf{N}}(\boldsymbol{0},\boldsymbol{I}_{n})$, , independent of $\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle$ so that:
$\boldsymbol{w}=\dfrac{(\boldsymbol{I}_{n}-\boldsymbol{u}\boldsymbol{u}^{{\sf T}})\boldsymbol{g}}{\|(\boldsymbol{I}_{n}-\boldsymbol{u}\boldsymbol{u}^{{\sf T}})\boldsymbol{g}\|}\,.$
For every $1\leq i\leq n$, we have
$|((\boldsymbol{I}_{n}-\boldsymbol{u}\boldsymbol{u}^{{\sf T}})\boldsymbol{g})_{i}|\leq|g_{i}|+\dfrac{|\langle\boldsymbol{u},\boldsymbol{g}\rangle|}{\sqrt{k}}\,.$
Since by assumption $k\log(n/k)\ll n$, with probability at least $1-C_{1}\exp(-k/2)$, $|\langle\boldsymbol{u},\boldsymbol{g}\rangle|\leq\sqrt{k\log(n/k)}$ and $\|(\boldsymbol{I}_{n}-\boldsymbol{u}\boldsymbol{u}^{{\sf T}})\boldsymbol{g}\|\geq\sqrt{n}/2$, so that
$i\in{\mathcal{L}}(\boldsymbol{g};C)\;\;\Rightarrow\;\;|w_{i}|\leq(2C+2)\cdot\sqrt{\frac{1}{n}\log(n/k)}\,.$
Therefore, by using Lemma 5.4 and $k\log(n/k)\gg n^{1/2}$, we get that with probability at least $1-C_{*}\exp(-n^{1/4})$, $|{\mathcal{A}}(\boldsymbol{w};C)|\leq\max\{\sqrt{k},k^{2}/n\}$ for some constant $C>0$, where
${\mathcal{A}}(\boldsymbol{w};C):=\left\{i:1\leq i\leq n,|w_{i}|\geq C\sqrt{\frac{1}{n}\log(n/k)}\right\}\,.$
By Lemma 5.3, we know that with probability at least $1-C_{*}\exp(-cn^{1/3})$ for some $C_{*},c>0$, $|\langle\boldsymbol{v}_{t_{0}(1+\delta)},\boldsymbol{u}\rangle|\geq c\delta^{2}$. Since $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|$ is stochastically increasing with $t$ (Fact F.2), we actually have that for any $t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}$, with the same probability $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|\geq c\delta^{2}$.
On the event $\mathcal{E}_{\delta}:=\{|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|\geq c\delta^{2}\}$, further suppose without loss of generality that $\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle>0$. As a consequence of the result above, if $i\in\text{supp}(\boldsymbol{u})$ and $i\notin{\mathcal{A}}(\boldsymbol{w};C)$, then
$u_{i}>0,i\notin{\mathcal{A}}(\boldsymbol{w};C)$ $\;\;\Rightarrow\;\;v_{t,i}\geq\frac{c\delta^{2}}{\sqrt{k}}-C\sqrt{\frac{1}{n}\log(n/k)}\,,$
$u_{i}<0,i\notin{\mathcal{A}}(\boldsymbol{w};C)$ $\;\;\Rightarrow\;\;v_{t,i}\leq-\frac{c\delta^{2}}{\sqrt{k}}+C\sqrt{\frac{1}{n}\log(n/k)}\,.$
Similarly, if $i\notin\text{supp}(\boldsymbol{u})$, then
$u_{i}=0,\;\;i\notin{\mathcal{A}}(\boldsymbol{w};C)\;\;\Rightarrow\;\;|v_{t,i}|\leq C\sqrt{\frac{1}{n}\log(n/k)}\,.$
We next choose $\delta=\delta_{n}$ such that $\sqrt{(k/n)\log(n/k)}\ll\delta_{n}\ll 1$. Hence, we conclude that
$i\in{\rm supp}(\boldsymbol{u})\setminus{\mathcal{A}}(\boldsymbol{w};C)\;\;\Rightarrow\;\;\hat{v}_{t_{i}}=\operatorname{sign}(\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle)\cdot u_{i}\,,$ (24)
whence
$|\langle\boldsymbol{u},\hat{\boldsymbol{v}}_{t}\rangle|\geq 1-\frac{2}{k}|{\mathcal{A}}(\boldsymbol{w};C)|=1-o_{n}(1)\,,$ (25)
with probability at least $1-C_{*}\exp(-n^{1/4}/3)$. On this event, by Eq. (23) we obtain that
$s_{t}\geq t\cdot(1-o_{n}(1))^{2}-4\cdot\sqrt{k\log(n/k)}\cdot\sqrt{t}$
For $t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}=(1+\delta)n/2$, we this implies that, for any fixed $\beta\in(0,1)$, with probability at least $1-C_{*}\exp(-n^{1/4}/3)$ (possibly adjusting the constant $C_{*}$).
Recalling that $\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)=\hat{\boldsymbol{m}}_{0}(\boldsymbol{y}_{t},t)\phi_{1}(\boldsymbol{y}_{t},t)$ for $t\in[t_{1},t_{2}]$, and $\|\hat{\boldsymbol{m}}_{0}(\boldsymbol{y},t)\|\leq 1$, we have, for $t_{1}=(1+\delta)t_{\mbox{\tiny\rm alg}}$, $t_{2}=n^{4}$ as defined above,
$J_{n}(t_{1},t_{2})=\int_{t_{1}}^{t_{2}}\mathbb{E}[\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\hat{\boldsymbol{m}}_{0}(\boldsymbol{y}_{t},t)\|^{2}]{\rm d}t\leq\int_{t_{1}}^{t_{2}}\mathbb{P}\left(\phi_{1}(\boldsymbol{y}_{t},t)=0\right)\leq C_{*}n^{4}\,e^{-cn^{1/3}}\,.$ (26)
Bounding $J_{n}(t_{2},\infty)$.
When $t\geq t_{2}$, we use a simple eigenvalue test. For a matrix $\boldsymbol{y}$, and time point $t$, we define $\phi_{2}(\boldsymbol{y},t)$ according to the following procedure:
  1. 1.
    Compute symmetrized matrix $\boldsymbol{A}=(1/2)(\boldsymbol{y}+\boldsymbol{y}^{{\sf T}})$.
  2. 2.
    Compute top eigenvalue $\lambda_{1}(\boldsymbol{A})$.
  3. 3.
    Return $1$ if $\lambda_{1}(\boldsymbol{A})\geq t/2$, and 00 otherwise.
Under the true model $\boldsymbol{y}_{t}=t\boldsymbol{x}-\boldsymbol{B}_{t}$, we have:
$\lambda_{1}(\boldsymbol{A}_{t})\geq t-\|\boldsymbol{W}_{t}\|_{{\mbox{\rm\tiny op}}}\geq t-t^{2/3}$
with error probability given by
$\mathbb{P}\left(\|\boldsymbol{W}_{t}\|_{{\mbox{\rm\tiny op}}}\geq t^{2/3}\right)\leq C\exp\left(-ct^{1/3}\right)\,,$
for constants $C,c>0$. Thus, we get that
$J_{n}(t_{2},\infty)$ $\leq 2\int_{t_{2}}^{\infty}\mathbb{P}(\|\boldsymbol{W}_{t}\|_{{\mbox{\rm\tiny op}}}\geq t^{2/3})\,{\rm d}t\leq C_{*}\int_{t_{2}}^{\infty}e^{-ct^{1/3}}{\rm d}t$
$\leq C_{**}e^{-ct_{2}^{1/3}}\leq C_{**}e^{-cn^{4/3}}\,.$ (27)
Claim M2 of Theorem 1 follows from Eqs. (22), (26), (27).
We finally consider claim M3. Equation (4) yields, for every $\ell\geq 1$:
$\hat{\boldsymbol{y}}_{\ell\Delta}$ $=\Delta\sum_{i=1}^{\ell}\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{(i-1)\Delta},(i-1)\Delta)+\sqrt{\Delta}\sum_{i=1}^{\ell}\boldsymbol{g}_{i\Delta}\,.$ (28)
We define
$\overline{\boldsymbol{m}}_{t_{1}}$ $:=\Delta\sum_{i=1}^{\ell_{1}}\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{(i-1)\Delta},(i-1)\Delta)\,,\,,\;\;\;\;\ell_{1}:=\lfloor t_{\mbox{\tiny\rm alg}}(1+\delta)/\Delta\rfloor\,.$ (29)
We further define the following auxiliary process for $t\geq\ell_{1}\Delta=\lfloor t_{1}/\Delta\rfloor\Delta$:
$\tilde{\boldsymbol{y}}_{t}=\overline{\boldsymbol{m}}_{t_{1}}+\boldsymbol{B}_{t}\,,\;\;\;\;$ (30)
where $(\boldsymbol{B}_{t}:\,t\geq 0)$ is a Brownian motion such that $\boldsymbol{B}_{j\Delta}=\sqrt{\Delta}\sum_{i=1}^{k}\boldsymbol{g}_{i\Delta}$. In particular, $\tilde{\boldsymbol{y}}_{\ell_{1}\Delta}=\hat{\boldsymbol{y}}_{\ell_{1}\Delta}$.
From triangle inequality, using Assumption 1, the condition $\|\hat{\boldsymbol{m}}_{0}(\boldsymbol{y},t)\|_{F}\leq 1$, and that by construction $\|\hat{\boldsymbol{m}}(\boldsymbol{y},t)\|_{F}\leq 1-\varepsilon$ for $t\leq(1-\gamma)t_{\mbox{\tiny\rm alg}}$, we get
$\|\overline{\boldsymbol{m}}\|_{F}\leq(1-\gamma)t_{\mbox{\tiny\rm alg}}(1-\varepsilon)+(\gamma+\delta)t_{\mbox{\tiny\rm alg}}\,.$
We claim that (with high probability) $\phi_{1}(\tilde{\boldsymbol{y}}_{t},t)=0$ simultaneously for all $t\in[t_{1},t_{2}]$. As a consequence, recalling the definition of $\hat{\boldsymbol{m}}$, we obtain that $\tilde{\boldsymbol{y}}_{t}=\hat{\boldsymbol{y}}_{t}$ for all $t\in[t_{1},t_{2}]$. In order to prove the claim, define, for $\ell\geq 1$:
$\mathcal{T}_{n}:=\big{\{}t^{+}_{\ell}=t_{1}+\dfrac{\ell-1}{n}:\ell\in{\mathbb{N}}\big{\}}\cap[t_{1},t_{2}]\,.$
For every $t\in[t_{1},t_{2}]$, consider the thresholded vector $\hat{\boldsymbol{v}}_{t}$ in the definition of thest $\phi_{1}$. We know that
$\langle\hat{\boldsymbol{v}}_{t},\tilde{\boldsymbol{y}}_{t}\hat{\boldsymbol{v}}_{t}\rangle=\langle\hat{\boldsymbol{v}}_{t},\overline{\boldsymbol{m}}\hat{\boldsymbol{v}}_{t}\rangle+\langle\hat{\boldsymbol{v}}_{t},\boldsymbol{B}_{t}\hat{\boldsymbol{v}}_{t}\rangle\leq(1-\gamma)t_{\mbox{\tiny\rm alg}}(1-\varepsilon)+(\gamma+\delta)t_{\mbox{\tiny\rm alg}}+\max_{\boldsymbol{v}\in\Omega_{n,k}}\langle\boldsymbol{v},\boldsymbol{B}_{t}\boldsymbol{v}\rangle\,.$ (31)
Using Lemma 5.1, we get that
$\mathbb{P}\left(\max_{\boldsymbol{v}\in\Omega_{n,k}}|\langle\boldsymbol{v},\boldsymbol{B}_{t}\boldsymbol{v}\rangle|\geq C\sqrt{\log{\binom{n}{k}}}\sqrt{t}\right)\leq 2\binom{n}{k}^{-C^{2}/2+2}\,.$
Note that $|\mathcal{T}_{n}|\leq n^{5}$, whence
$\mathbb{P}\left(\exists t\in\mathcal{T}_{n}:\max_{\boldsymbol{v}\in\Omega_{n,k}}|\langle\boldsymbol{v},\boldsymbol{W}_{t_{\ell}}\boldsymbol{v}\rangle|\geq C\sqrt{\log{\binom{n}{k}}}\sqrt{t_{\ell}}\right)\leq 2\binom{n}{k}^{-C^{2}/2+2}n^{5}\,.$
For $t\in[t_{\ell},t_{\ell+1}]$, we have
$\max_{t_{\ell}\leq t\leq t_{\ell+1}}\left\{\max_{\boldsymbol{v}\in\Omega_{n,k}}|\langle\boldsymbol{v},\boldsymbol{W}_{t}\boldsymbol{v}\rangle|-\max_{\boldsymbol{v}\in\Omega_{n,k}}|\langle\boldsymbol{v},\boldsymbol{W}_{t_{\ell}}\boldsymbol{v}\rangle|\right\}\leq\max_{t_{\ell}\leq t\leq t_{\ell+1}}\|\boldsymbol{W}_{t}-\boldsymbol{W}_{t_{\ell}}\|_{{\mbox{\rm\tiny op}}}\,.$
Using Lemma 6.3, we get that $\max_{t_{\ell}\leq t\leq t_{\ell+1}}\|\boldsymbol{W}_{t}-\boldsymbol{W}_{t_{\ell}}\|_{{\mbox{\rm\tiny op}}}\leq 16\sqrt{(t_{\ell+1}-t_{\ell})n}=16$ with probability at least $1-2\exp(-32n)$. Taking a union bound over $\mathcal{T}_{n}$, we get that
$\mathbb{P}\left(\exists t\in[t_{1},t_{2}]:\max_{\boldsymbol{v}\in\Omega_{n,k}}|\langle\boldsymbol{v},\boldsymbol{W}_{t}\boldsymbol{v}\rangle|\geq C\sqrt{\log\binom{n}{k}}\sqrt{t}\right)\leq 2\binom{n}{k}^{-C^{2}/2+2}n^{5}+2n^{5}\,e^{-32n}\,,$
for possibly a different constant $C>0$.
Using Eq. (31) and the last estimate, we obtain that
$\mathbb{P}\left(\exists t\in[t_{1},t_{2}]:\langle\hat{\boldsymbol{v}}_{t},\tilde{\boldsymbol{y}}_{t}\hat{\boldsymbol{v}}_{t}\rangle\geq b_{n}t_{\mbox{\tiny\rm alg}}+C\sqrt{\log\binom{n}{k}}\sqrt{t}\right)\leq 2\binom{n}{k}^{-C^{2}/2+2}n^{5}+2n^{5}\exp(-32n)\,,$ (32)
where $b_{n}:=(1-\gamma)(1-\varepsilon)+(\gamma+\delta_{n})$. Since $\delta_{n}=o_{n}(1)$, we have $b_{n}\to(1-\gamma)(1-\varepsilon)+\gamma<1$. Hence, we get that for $t\geq t_{1}$, and all $n$ large enough
$b_{n}t_{\mbox{\tiny\rm alg}}+C\sqrt{\log\binom{n}{k}}\sqrt{t}<b_{*}t$
for some constant $b_{*}\in(0,1)$ large enough. Therefore Eq. (32) implies that $\phi_{1}(\tilde{\boldsymbol{y}}_{t},t)=0$ simultaneously for all $t\in[t_{1},t_{2}]$, with high probability. We conclude that, with high probability $\hat{\boldsymbol{y}}_{t}=\tilde{\boldsymbol{y}}_{t}$ throughout $t\in[t_{1},t_{2}]$.
Finally, we extend the analysis to $t\in[t_{2},\infty)$ by proving that, with high probability, $\phi_{2}(\tilde{\boldsymbol{y}}_{t},t)=0$ and hence $\hat{\boldsymbol{y}}_{t}=\tilde{\boldsymbol{y}}_{t}$ for all $t\in[t_{2},\infty)$. We use (for $\boldsymbol{A}_{t}=(\tilde{\boldsymbol{y}}_{t}+\tilde{\boldsymbol{y}}_{t}^{{\sf T}})/2$, $\boldsymbol{W}_{t}=(\boldsymbol{B}_{t}+\boldsymbol{B}_{t}^{{\sf T}})/2$)
$\lambda_{1}(\boldsymbol{A}_{t})\leq(1-\gamma)t_{\mbox{\tiny\rm alg}}(1-\varepsilon)+(\delta+\gamma)t_{\mbox{\tiny\rm alg}}+\lambda_{1}(\boldsymbol{W}_{t})$ (33)
Following exactly the argument as in the proof of Theorem 2 (in particular, Subsection 6.3.2), we get that $\lambda_{1}(\boldsymbol{W}_{t})\leq t/3$ simultaneously for all $t\geq t_{2}$, with high probability. On this event, $\lambda_{1}(\boldsymbol{A}_{t})<t/2$ with high probability (because $t/6\gg(1-\gamma)t_{\mbox{\tiny\rm alg}}(1-\varepsilon)+(\gamma+\delta)t_{\mbox{\tiny\rm alg}}$). Hence, with high probability $\phi_{2}(\tilde{\boldsymbol{y}}_{t},t)=0$ and hence $\hat{\boldsymbol{y}}_{t}=\tilde{\boldsymbol{y}}_{t}$ for all $t\in[t_{2},\infty)$.
We thus proved that, with high probability, $\hat{\boldsymbol{y}}_{t}=\tilde{\boldsymbol{y}}_{t}$ for all $t\geq t_{1}=(1+\delta)t_{\mbox{\tiny\rm alg}}$, whence $\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{t},t)=0$ as well (because $\phi_{1}(\tilde{\boldsymbol{y}}_{t},t)=0$ for $t\in[t_{1},t_{2}]$ and $\phi_{2}(\tilde{\boldsymbol{y}}_{t},t)=0$ for $t\in[t_{2},\infty)$). Claim M3 thus follows.
6 Proof of Theorem 2
6.1 Properties of the estimator $\hat{\boldsymbol{m}}(\,\cdot\,)$
Proposition 6.1.
Assume $\sqrt{n}\ll k\ll n$, and note that in this case $t_{\mbox{\tiny\rm alg}}(n,k)=n/2$. Let $\hat{\boldsymbol{m}}(\,\cdot\,)$ be the estimator of Algorithm 1 with input parameter $\varepsilon$. For every $\delta>0$, there exists $\varepsilon>0$ such that
$\sup_{t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\left[\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\right]\leq C\,e^{-n\varepsilon^{2}/64k}\,.$ (34)
The proof of this proposition is standard, and will be presented in Appendix F. We note that the rate in Equation (34) gets slower the closer $k$ is to $n$; it is super-polynomial if $n\gg k\log n$.
By definition, when $\sqrt{n}\ll k\ll n$ and $t<t_{\mbox{\tiny\rm alg}}(n,k)$, the Algorithm 1 returns $\hat{\boldsymbol{m}}(\boldsymbol{y},t)=\boldsymbol{0}$, so we automatically have the following result.
Proposition 6.2.
For any fixed $\delta>0$, and $t\leq(1-\delta)t_{\mbox{\tiny\rm alg}}$, we have $\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|=1$.
6.2 Auxiliary lemmas
The following lemmas are needed for the analysis of the generated diffusion. Their proofs are deferred to Appendices G, H, I, J.
Lemma 6.3.
Let $\boldsymbol{W}_{t}$ be a ${\sf GOE}$ process. Then for each time $t_{0}\geq 0$,
$\mathbb{P}\left(\max_{0\leq t\leq t_{0}}\|\boldsymbol{W}_{t}\|_{{\mbox{\rm\tiny op}}}\geq 16\sqrt{t_{0}n}\right)\leq 2\exp\left(-32n\right).$
Lemma 6.4.
Let $\boldsymbol{W}_{t}$ be a ${\sf GOE}$ process, and let $\boldsymbol{v}_{t}$ be any eigenvector of $\boldsymbol{W}_{t}$ for every $t\geq 0$. Define the set
$A(\boldsymbol{v}_{t};C)=\left\{i:1\leq i\leq n,|v_{ti}|\geq\dfrac{C\sqrt{\log(n/k)}}{\sqrt{n}}\right\}$
Then for any $C>4$, we have
$\mathbb{P}\left(|A(\boldsymbol{v}_{t};C)|\geq\max\{\sqrt{k},k^{2}/n\}\right)=O\left(\exp(-(1/3)n^{1/4})\right)$
As a consequence, using this eigenvector, $\hat{\boldsymbol{m}}$ will evaluate to $\boldsymbol{0}$ per line 8 of Algorithm 1.
Lemma 6.5.
Let $\boldsymbol{W}_{t}$ be a ${\sf GOE}$ process, and for each $t$, let $\boldsymbol{v}_{t}$ be a top eigenvector of $\boldsymbol{W}_{t}$. Then for any times $t_{0}\leq t_{1}$, with probability at least $1-2\exp(-32n)$,
$\sup_{t_{0}\leq t\leq t_{1}}|\langle\boldsymbol{v}_{t},\boldsymbol{W}_{t_{0}}\boldsymbol{v}_{t}\rangle-\lambda_{1}(\boldsymbol{W}_{t_{0}})|\leq 16\sqrt{n(t_{1}-t_{0})}.$
Lemma 6.6 (Concentration for deformed GOE model).
Consider the model $\boldsymbol{Y}=\theta\boldsymbol{v}\boldsymbol{v}^{{\sf T}}+\boldsymbol{W}$ for $\boldsymbol{W}\sim{\sf GOE}(n)/\sqrt{n}$ and $\theta>1$ a constant, $\boldsymbol{v}$ a unit vector. Let $\boldsymbol{v}_{1}(\boldsymbol{Y})$ be the top eigenvector of $\boldsymbol{Y}$. Define $(x^{\star},u^{\star})=(\theta+1/\theta,1-1/\theta^{2})$. For any closed set $F$ such that $d((x^{\star},u^{\star}),F)>0$, there exists a constant $c>0$ such that
$\mathbb{P}\left((\lambda_{1}(\boldsymbol{Y}),\langle\boldsymbol{v}_{1}(\boldsymbol{Y}),\boldsymbol{v}\rangle^{2})\in F\right)\leq\exp(-cn)$
for all $n$ large enough.
Remark 6.1.
We only use Lemma 6.6 for the alignment $\langle\boldsymbol{v}_{1}(\boldsymbol{Y}),\boldsymbol{v}\rangle^{2}$.
6.3 Analysis of the diffusion process: Proof of Theorem 2
We will prove Theorem 2 for $1\gg\varepsilon\geq C\sqrt{\log(n/k)/(n/k)}$ for some sufficiently large constant $C$.
Suppose that we generate the following diffusion, with $(\boldsymbol{z}_{t})_{t\geq 0}$ a standard $n^{2}$-dimensional Brownian motion, and $\hat{\boldsymbol{y}}_{0}=\boldsymbol{0}$:
$\hat{\boldsymbol{y}}_{\ell\Delta}=\hat{\boldsymbol{y}}_{(\ell-1)\Delta}+\Delta\cdot\hat{\boldsymbol{m}}\left(\hat{\boldsymbol{y}}_{(\ell-1)\Delta},(\ell-1)\Delta\right)+\left(\boldsymbol{z}_{\ell\Delta}-\boldsymbol{z}_{(\ell-1)\Delta}\right).$
We will prove that the generated diffusion never passes the termination conditions (c.f. Algorithm 1, lines 3, 5, 8).
6.3.1 Analysis up to an intermediate time
Define $t_{\mbox{\tiny\rm between}}=n^{2}$. Following the same strategy with Section A.2, we will first show that $\hat{\boldsymbol{m}}=\boldsymbol{0}$ up to $t_{\mbox{\tiny\rm between}}$ with high probability by analyzing only the noise process (in short, if $\hat{\boldsymbol{m}}=\boldsymbol{0}$ always, our generated diffusion coincides with the noise process). Our strategy is of the same nature as that of Section A.2. Indeed, we will attempt to prove that $\hat{\boldsymbol{m}}=\boldsymbol{0}$ simultaneously for all $t$, with high probability. In this phase ($0\leq t\leq t_{\mbox{\tiny\rm between}}$), we will show that $|\hat{S}|<k/2$ (c.f. definition in Algorithm 1, lines 7, 8) for $0\leq t\leq t_{\mbox{\tiny\rm between}}$, with high probability ($\boldsymbol{v}_{t}$ is the top eigenvector of $\boldsymbol{A}_{t}$, c.f. Algorithm 1). Note that line 5 of Algorithm 1 is not relevant in this phase. We first show this for a sequence of time points $\{t_{\ell}\}_{\ell\geq 1}$, then control the in-between fluctuations. We can set $t_{1}$ to be any value in $[0,n/2)$, as the algorithm returns 00 if $t<t_{\mbox{\tiny\rm alg}}=n/2$ anyway. We denote the ${\sf GOE}$ process
$\boldsymbol{W}_{t}=\dfrac{\boldsymbol{B}_{t}+\boldsymbol{B}_{t}^{{\sf T}}}{2}=\sqrt{t}\boldsymbol{A}_{t}.$
It is clear that the eigenvectors of $\boldsymbol{W}_{t}$ and $\boldsymbol{A}_{t}$ coincide.
We choose the following time points:
$t_{\ell}=\dfrac{n}{2}-1+\dfrac{\ell}{n^{4}}.$
To exceed $t_{\mbox{\tiny\rm between}}=n^{2}$, we will need $n^{6}$ values of $\ell$. By union bound from Lemma 6.4 (recall also the definition of the set $A(\boldsymbol{v};C)$ from this Lemma),
$\mathbb{P}\left(\exists 1\leq\ell\leq n^{6}:|A(\boldsymbol{v}_{t_{\ell}};C)|\geq\max\{\sqrt{k},k^{2}/n\}\right)\leq O\left(\exp(-(1/3)n^{1/4}+6\log n)\right)$ (35)
Next, we will control the in-between fluctuations; specifically, we would like to show that $\max_{t_{\ell}\leq t\leq t_{\ell+1}}|A(\boldsymbol{v}_{t};C)|\leq C_{0}\max\{\sqrt{k},k^{2}/n\}$ simultaneously for many values of $\ell$ (with high probability), for some constant $C_{0}>0$. Our approach is as follows.
  • (i)
    Let $\boldsymbol{v}_{t}$ be a top eigenvector of $\boldsymbol{W}_{t}$. If $t$ is close to $t_{\ell}$, then $\boldsymbol{v}_{t}$ is an approximate solution to the equation (in $\boldsymbol{v}$):
    $\boldsymbol{v}^{{\sf T}}\boldsymbol{W}_{t_{\ell}}\boldsymbol{v}=\lambda_{1}(\boldsymbol{W}_{t_{\ell}})$
  • (ii)
    $\boldsymbol{v}_{t}$ can be written in the coordinate system of the orthonormal eigenvectors $\boldsymbol{U}_{t_{\ell}}=[\boldsymbol{u}_{1}|\cdots|\boldsymbol{u}_{n}]$ of $\boldsymbol{W}_{t_{\ell}}$, corresponding to decreasing eigenvalues $\lambda_{1}(\boldsymbol{W}_{t_{\ell}})\geq\cdots\geq\lambda_{n}(\boldsymbol{W}_{t_{\ell}})$. Namely, $\boldsymbol{v}_{t}=\boldsymbol{U}_{t_{\ell}}\boldsymbol{U}_{t_{\ell}}^{{\sf T}}\boldsymbol{v}_{t}=\boldsymbol{U}_{t_{\ell}}\boldsymbol{w}$ with $\|\boldsymbol{w}\|=1$.
Define $p_{n}=P\left(|\lambda_{1}(\boldsymbol{W}_{1})-\lambda_{7}(\boldsymbol{W}_{1})|\leq n^{-C^{\prime}-1/2}\right)$ for any $C^{\prime}>0$ (here we take $m=7$). We use the following result (we have accounted for the scaling).
Lemma 6.7 (Corollary 2.5, [NTV17]).
Let $\boldsymbol{W}_{1}\sim(1/\sqrt{2}){\sf GOE}(n)$. For any fixed $l\geq 1,C^{\prime}>0$, there exists a constant $c_{0}=c_{0}(l,C^{\prime})$ such that
$\mathbb{P}\left(\lambda_{1}(\boldsymbol{W}_{1})-\lambda_{1+l}(\boldsymbol{W}_{1})\leq\dfrac{1}{2}n^{-C^{\prime}-1/2}\right)\leq c_{0}n^{-C^{\prime}\cdot\frac{l^{2}+2l}{3}}.$
We materialize our approach above. We can write, with $\boldsymbol{v}_{t}=\boldsymbol{U}_{t_{\ell}}\boldsymbol{w}$:
$\boldsymbol{v}_{t}^{{\sf T}}\boldsymbol{W}_{t_{\ell}}\boldsymbol{v}_{t}=\boldsymbol{w}^{{\sf T}}\boldsymbol{D}_{t_{\ell}}\boldsymbol{w}=\sum_{i=1}^{n}(D_{t_{\ell}})_{ii}w_{i}^{2}.$
We then obtain that
$\boldsymbol{v}_{t}^{{\sf T}}\boldsymbol{W}_{t_{\ell}}\boldsymbol{v}_{t}-\lambda_{1}(\boldsymbol{W}_{t_{\ell}})\leq\sum_{i=8}^{n}(\lambda_{i}(\boldsymbol{W}_{t_{\ell}})-\lambda_{1}(\boldsymbol{W}_{t_{\ell}}))w_{i}^{2}<-\dfrac{1}{2}\sqrt{t_{\ell}}n^{-3/2}\sum_{i=8}^{n}w_{i}^{2}$
with probability at least $1-p_{n}\geq 1-c_{0}n^{-8}$, from Lemma 6.7 and $\boldsymbol{W}_{t_{\ell}}\sim\sqrt{t_{\ell}}\boldsymbol{W}_{1}$. Now from Lemma 6.5, we know that with probability at least $1-2\exp(-32n)$,
$\boldsymbol{v}_{t}^{{\sf T}}\boldsymbol{W}_{t_{\ell}}\boldsymbol{v}_{t}-\lambda_{1}(\boldsymbol{W}_{t_{\ell}})\geq-16\sqrt{n(t_{\ell+1}-t_{\ell})}.$
With probability at least $1-c_{0}n^{-8}-2\exp(-32n)$, both of these statements are true, uniformly over $t_{\ell}\leq t\leq t_{\ell+1}$, leading to
$32\sqrt{\dfrac{t_{\ell+1}-t_{\ell}}{t_{\ell}}}\geq n^{-3/2}\sum_{i=8}^{n}w_{i}^{2}.$
A simple bit of algebra shows that
$\sqrt{\dfrac{t_{\ell+1}-t_{\ell}}{t_{\ell}}}\leq 2n^{-5/2}\Rightarrow\sum_{i=8}^{n}w_{i}^{2}\leq 64n^{-1}.$
Consider the first $7$ eigenvectors $\{\boldsymbol{u}_{t_{\ell},i}\}_{i=1}^{7}$ of $\boldsymbol{W}_{t_{\ell}}$. Let
$A=\bigcup_{i=1}^{7}A(\boldsymbol{u}_{t_{\ell},i};C).$
From Lemma 6.4 and a union bound, that $|A|\leq 7\max\{\sqrt{k},k^{2}/n\}$ with probability at least $1-O(\exp(-(1/3)n^{1/4}))$. For every $j\in A^{c}$, we have
$|v_{tj}|\leq\sum_{i=1}^{n}|w_{i}|\cdot|u_{t_{\ell},i,j}|<\sum_{i=1}^{7}|w_{i}|\cdot\dfrac{C\sqrt{\log(n/k)}}{\sqrt{n}}+\sum_{i=8}^{n}|w_{i}|\leq\dfrac{C^{\prime}\sqrt{\log(n/k)}}{\sqrt{n}}+\dfrac{64}{\sqrt{n}}<C^{\prime\prime}\sqrt{\dfrac{\log(n/k)}{n}}$
for a large enough constant $C^{\prime\prime}>0$. This means that with probability at least $1-O(n^{-8})$,
$\sup_{t_{\ell}\leq t\leq t_{\ell+1}}|A(\boldsymbol{v}_{t};C^{\prime\prime})|\leq 7\max\{\sqrt{k},k^{2}/n\}<k/2$
From Equation (35) and a union bound over $\ell\geq 1$, we know that with high probability,
$\sup_{n/2-1\leq t\leq n^{2}}|A(\boldsymbol{v}_{t};C)|\leq 7\max\{\sqrt{k},k^{2}/n\}$
for some absolute constant $C>0$, meaning that $\hat{\boldsymbol{m}}=0$ up to $t_{\mbox{\tiny\rm between}}$, as long as
$\varepsilon>\dfrac{C\sqrt{\log(n/k)}}{\sqrt{n/k}}$
6.3.2 Analysis to the infinite horizon
We will prove that simultaneously for all $t\geq n^{2}$, Algorithm 1 always terminates at line 5, or that $\lambda_{1}(\boldsymbol{W}_{t})\leq t/2$. Similar to Subsection 6.3.1, we choose the following sequence of time points for all $\ell\geq 1$:
$t_{\ell}^{(2)}=n^{2}+\ell-1$
By standard Gaussian concentration and the Bai-Yin theorem, we get, for instance, the following tail bound (constants are loose) for all $x\geq 0$:
$\mathbb{P}\left(\lambda_{1}\left(\dfrac{\boldsymbol{W}_{t}}{\sqrt{t}}\right)\geq 4\sqrt{n}+x\right)\leq 2\exp\left(-\dfrac{x^{2}}{2}\right)$
Set $x=\dfrac{\sqrt{t}}{8}$. Since $t\geq n^{2}$, we have $x\geq n/8$, and so $x+4\sqrt{n}\leq 2x$ for $n$ large enough. Consequently,
$\mathbb{P}\left(\lambda_{1}\left(\dfrac{\boldsymbol{W}_{t}}{\sqrt{t}}\right)\geq\dfrac{\sqrt{t}}{4}\right)\leq 2\exp\left(-c_{1}t\right)$
for some universal constant $c_{1}>0$. A union bound for the chosen points gives:
$\mathbb{P}\left(\exists\ell\geq 1:\lambda_{1}\left(\boldsymbol{W}_{t_{\ell}^{(2)}}\right)\geq t_{\ell}^{(2)}/4\right)\leq 2\sum_{\ell=1}^{\infty}\exp\left(-c_{1}t_{\ell}^{(2)}\right)=2\exp(-c_{1}n^{2})\sum_{\ell=1}^{\infty}\exp(-c_{1}(\ell-1))\lesssim\exp(-c_{1}n^{2})$ (36)
Next we control the in-between fluctuations. From a simple modification of Lemma 6.3, we have
$\mathbb{P}\left(\sup_{t_{\ell}^{(2)}\leq t\leq t_{\ell+1}^{(2)}}\left|\lambda_{1}\left(\boldsymbol{W}_{t_{\ell}^{(2)}}\right)-\lambda_{1}\left(\boldsymbol{W}_{t}\right)\right|\geq 16\sqrt{(t_{\ell+1}^{(2)}-t_{\ell}^{(2)})\cdot t_{\ell}^{(2)}n}\right)\leq 2\exp\left(-32nt_{\ell}^{(2)}\right)$
so that by union bound
$\mathbb{P}\left(\exists\ell\geq 1:\sup_{t_{\ell}^{(2)}\leq t\leq t_{\ell+1}^{(2)}}\left|\lambda_{1}\left(\boldsymbol{W}_{t_{\ell}^{(2)}}\right)-\lambda_{1}\left(\boldsymbol{W}_{t}\right)\right|\geq 16\sqrt{(t_{\ell+1}^{(2)}-t_{\ell}^{(2)})\cdot t_{\ell}^{(2)}n}\right)\lesssim\exp(-32n^{3})$ (37)
Consider the intersection of events described in Equations (36) and (37):
$A=\left\{\exists\ell\geq 1:\lambda_{1}\left(\boldsymbol{W}_{t_{\ell}^{(2)}}\right)\geq t_{\ell}^{(2)}/4\right\}\cup\left\{\exists\ell\geq 1:\sup_{t_{\ell}^{(2)}\leq t\leq t_{\ell+1}^{(2)}}\left|\lambda_{1}\left(\boldsymbol{W}_{t_{\ell}^{(2)}}\right)-\lambda_{1}\left(\boldsymbol{W}_{t}\right)\right|\geq 16\sqrt{(t_{\ell+1}^{(2)}-t_{\ell}^{(2)})\cdot t_{\ell}^{(2)}n}\right\}$
For each $t\geq n^{2}$, let $t_{\ell}$ be largest such that $t_{\ell}\leq t<t_{\ell+1}$. On $A$, we have
$\lambda_{1}(\boldsymbol{W}_{t})\leq\dfrac{t_{\ell}^{(2)}}{4}+16\sqrt{(t_{\ell+1}^{(2)}-t_{\ell}^{(2)})\cdot t_{\ell}^{(2)}n}=\dfrac{t_{\ell}^{(2)}}{4}+16\sqrt{t_{\ell}^{(2)}n}\leq\dfrac{t_{\ell}^{(2)}}{2}\leq\dfrac{t}{2}$
for $n$ large enough, since $t_{\ell}^{(2)}\geq n^{2}\gg n$. Hence the algorithm always returns $\boldsymbol{0}$ with high probability, and we are done.
Acknowledgements
This work was supported by the NSF through award DMS-2031883, the Simons Foundation through Award 814639 for the Collaboration on the Theoretical Foundations of Deep Learning, and the ONR grant N00014-18-1-2729.
Appendix A Proof of Theorem 3
A.1 Properties of the estimator $\hat{\boldsymbol{m}}(\,\cdot\,)$
Proposition A.1.
Assume $(\log n)^{2}\ll k\ll\sqrt{n}$ and let $\hat{\boldsymbol{m}}(\,\cdot\,)$ be the estimator defined in Algorithm 2. Recall that in this case, $t_{\mbox{\tiny\rm alg}}=k^{2}\log(n/k^{2})$. Then for any $\delta>0$ there exists $\varepsilon>0$ such that, letting $s=\sqrt{(1+\varepsilon)\log(n/k^{2})}$, we have
$\sup_{t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{P}(\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)\neq\boldsymbol{x})=O(n^{-D})$ (38)
for any fixed $D>0$.
The proof of this proposition is a modification of the one in [CLR17], and will be presented in Appendix C. Note that Proposition A.1 directly implies the first inequality in Condition M2v of Theorem 3, as $\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|\leq 2$.
By definition, when $t<t_{\mbox{\tiny\rm alg}}$, the algorithm will return $\hat{\boldsymbol{m}}=\boldsymbol{0}$, so we automatically have the following, which implies the second inequality in Condition M2v.
Proposition A.2.
For any fixed $\delta>0$, and $t\leq(1-\delta)t_{\mbox{\tiny\rm alg}}$, we have $\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|=1$.
In the proof of Theorem 3, we will also make use of the following estimates, whose proof is deferred to the Appendix D.
Lemma A.3.
Let $(\boldsymbol{w}_{t}:t\geq 0)$ be a process defined as
$\boldsymbol{w}_{t}=\dfrac{1}{2}\left\{(\boldsymbol{B}_{t}+\sqrt{\varepsilon}\boldsymbol{g}_{t})+(\boldsymbol{B}_{t}+\sqrt{\varepsilon}\boldsymbol{g}_{t})^{{\sf T}}\right\}$
where $\boldsymbol{B},\boldsymbol{g}$ are independent $n^{2}$-dimensional Brownian motions, and $10\leq\varepsilon<$ 1. Then, for any $0\leq t_{0}\leq t_{1}$, and $t\geq 0$, $s\geq 1$,
$\mathbb{P}\Big{(}\max_{t_{0}\leq t\leq t_{1}}\|\boldsymbol{w}_{t}-\boldsymbol{w}_{t_{0}}\|_{F}\geq 4\sqrt{(t_{1}-t_{0})s}\cdot n\Big{)}$ $\leq 2\,e^{-n^{2}s/4}\,,$ (39)
$\mathbb{P}\big{(}\|\boldsymbol{w}_{t}\|_{F}\geq 4\sqrt{ts}\cdot n\big{)}$ $\leq 2\,e^{-n^{2}s/4}\,.$ (40)
A.2 Analysis of the diffusion process: Proof of Theorem 3
We are left to prove that Condition M3v of Theorem 3 holds.
For that purpose, we make the following choices about Algorithm 2:
  • (C1)
    We select the constants in the algorithm to be $\varepsilon_{n}=o_{n}(1)$ and $s_{n}=\sqrt{(1+\varepsilon_{n})\log(n/k^{2})}$. We will use the shorthands $s=s_{n}$ and $\varepsilon=\varepsilon_{n}$, unless there is ambiguity.
  • (C2)
    The process $(\boldsymbol{g}_{t})_{t\geq 0}$ used in Algorithm 2 follows a $n^{2}$-dimensional Brownian motion.
Note that Propositions A.1, A.2 hold under these choices, and in particular $\boldsymbol{g}_{t}\sim{\mathsf{N}}({\boldsymbol{0}},t{\boldsymbol{I}}_{n\times n})$ at all times. Also the sequence of random vectors $\boldsymbol{g}_{\ell\Delta}$, $\ell\in{\mathbb{N}}$ can be generated via $\boldsymbol{g}_{\ell\Delta}=\boldsymbol{g}_{(\ell-1)\Delta}+\sqrt{\Delta}\hat{\boldsymbol{g}}_{\ell}$, for some i.i.d. standard normal vectors $\{\hat{\boldsymbol{g}}_{\ell}\}_{\ell\geq 0}$.
Letting $(\boldsymbol{z}_{t})_{t\geq 0}$ a standard Brownian motion in ${\mathbb{R}}^{n\times n}$, and $\hat{\boldsymbol{y}}_{0}=\boldsymbol{0}$ we can rewrite the approximate diffusion (4) as follows (for $t\in{\mathbb{N}}\cdot\Delta$)
$\hat{\boldsymbol{y}}_{t+\Delta}=\hat{\boldsymbol{y}}_{t}+\Delta\cdot\hat{\boldsymbol{m}}\left(\hat{\boldsymbol{y}}_{t},t\right)+\left(\boldsymbol{z}_{t+\Delta}-\boldsymbol{z}_{t}\right)\,.$ (41)
We further define
$\boldsymbol{w}_{t}=\dfrac{1}{2}\big{\{}(\boldsymbol{z}_{t}+\sqrt{\varepsilon}\boldsymbol{g}_{t})+(\boldsymbol{z}_{t}+\sqrt{\varepsilon}\boldsymbol{g}_{t})^{{\sf T}}\big{\}}\,.$ (42)
It is easy to see that $(c(\varepsilon)\boldsymbol{w}_{t}:t\geq 0)$ is a ${\sf GOE}(n)$ process for $c(\varepsilon):=((1+\varepsilon)/2)^{-1/2}$. The key technical estimate in the proof of Theorem 3, Condition M3v is stated in the next proposition.
Proposition A.4.
Let $(\boldsymbol{w}_{t}:t\geq 0)$ be defined as per Eq. (42), and assume $\varepsilon_{n}=o_{n}(1)$ and $s_{n}=\sqrt{(1+\varepsilon_{n})\log(n/k^{2})}$. Further, assume $k\geq C(\log n)^{5/2}$ for some sufficiently large absolute constant $C>0$. Then
$\lim_{n\to\infty}\mathbb{P}\left\{\|\eta_{s}(\boldsymbol{w}_{t}/\sqrt{t})\|_{{\mbox{\rm\tiny op}}}\leq k+\sqrt{t}/s\;\forall t\geq 1\right\}=1\,.$ (43)
Before proving this proposition, let us show that it implies Condition M3v of Theorem 3. Indeed we claim that, with high probability, for all $\ell\in{\mathbb{N}}$, $\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{\ell\Delta},\ell\Delta)={\boldsymbol{0}}$ and $\hat{\boldsymbol{y}}_{\ell\Delta}=\boldsymbol{z}_{\ell{\Delta}}$. This is proven by induction over $\ell$. Indeed, if it holds up to a certain $\ell-1\in{\mathbb{N}}$, then we have $\hat{\boldsymbol{y}}_{\ell\Delta}=\boldsymbol{z}_{\ell\Delta}$ by Eq. (41) whence it follows that $\boldsymbol{A}_{t,+}=\boldsymbol{w}_{t}/\sqrt{t}$, for $t=\ell\Delta$ (c.f. Algorithm 2, line 4) and therefore $\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{t},t)={\boldsymbol{0}}$ by Proposition A.4 (because the condition in Algorithm 2, line 6, is never passed).
We therefore have
$\inf_{\ell\geq 0}W_{1}(\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{\ell\Delta},\ell\Delta),\boldsymbol{x})\geq\inf_{\ell\geq 0}\mathbb{P}(\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{\ell\Delta},\ell\Delta)={\boldsymbol{0}})=1-o_{n}(1)\,.$ (44)
This concludes the proof of Theorem 3. We next turn to proving Proposition A.4.
Proof of Proposition A.4.
We follow a strategy analogous to the proof of the Law of Iterated Logarithm. We choose a sparse sequence of time points $\{t_{\ell}\}_{\ell=1}^{\infty}$, and $(i)$ establish the statement jointly for these time points, and $(ii)$ control deviations in between. In particular, we consider
$t_{\ell}=\left(1+\dfrac{\ell-1}{n^{3}}\right)^{2}$
for all $\ell\geq 1$.
We first show that simultaneously for all $\ell\geq 1$, we have $\max_{i,j}\left|(\boldsymbol{w}_{t_{\ell}})_{ij}/\sqrt{t_{\ell}}\right|\leq 8t_{\ell}^{1/4}\sqrt{\log n}$. We have, by sub-gaussianity of $(\boldsymbol{w}_{t_{\ell}})_{ij}$ and a union bound (here we account also for the case where $i=j$, in which there is an inflated variance), along with $\varepsilon=o_{n}(1)$: using the bound $2xy\geq x+y$ when $x,y\geq 1$ for $x=t_{\ell}^{1/2},y=\log n$. Taking a union bound once again over $\ell$, we have
$\mathbb{P}\left(\exists\ell\geq 1,1\leq i,j\leq n:|(\boldsymbol{w}_{t_{\ell}})_{ij}|\geq 8t_{\ell}^{3/4}\sqrt{\log n}\right)\leq n^{-6}\cdot\sum_{\ell=0}^{\infty}\exp\left(-8\cdot\left(1+\dfrac{\ell}{n^{3}}\right)\right)$
We have, as the summands form a decreasing function of $\ell$ integer:
$\sum_{\ell=0}^{\infty}\exp\left(-8\cdot\left(1+\dfrac{\ell}{n^{3}}\right)\right)\leq$ $C+\int_{0}^{\infty}\exp\left(-\dfrac{8x}{n^{3}}\right){\rm d}x\leq Cn^{3}\,.$ (45)
We thus obtain that
$\mathbb{P}\left(\exists\ell\geq 1,1\leq i,j\leq n:|(\boldsymbol{w}_{t_{\ell}})_{ij}|\geq 8t_{\ell}^{3/4}\sqrt{\log n}\right)=O(n^{-3})\,.$ (46)
The point of this calculation is that simultaneously for all $\ell\geq 1$, we can truncate the entries of $\eta_{s}(\boldsymbol{w}_{t_{\ell}}/\sqrt{t_{\ell}})$ by $8t_{\ell}^{1/4}\sqrt{\log n}$ without worry.
Namely, for each $\ell\geq 1$, $\vartheta_{\ell}=8t_{\ell}^{1/4}\sqrt{\log n}$ we define $\tilde{\boldsymbol{w}}_{t_{\ell}}\in{\mathbb{R}}^{n\times n}$ by
$(\tilde{\boldsymbol{w}}_{t_{\ell}})_{ij}:=\begin{cases}\eta_{s}(\boldsymbol{w}_{t_{\ell}}/\sqrt{t_{\ell}})&\mbox{ if }\;\;|\eta_{s}(\boldsymbol{w}_{t_{\ell}}/\sqrt{t_{\ell}})|\leq\vartheta_{\ell}\,,\\ \vartheta_{\ell}&\mbox{ if }\;\;\eta_{s}(\boldsymbol{w}_{t_{\ell}}/\sqrt{t_{\ell}})>\vartheta_{\ell}\,,\\ -\vartheta_{\ell}&\mbox{ if }\eta_{s}(\boldsymbol{w}_{t_{\ell}}/\sqrt{t_{\ell}})<-\vartheta_{\ell}\,.\end{cases}$ (47)
By Eq. (46), we have
$\mathbb{P}\left(\exists\ell\geq 1\;:\eta_{s}(\boldsymbol{w}_{t_{\ell}}/\sqrt{t_{\ell}})\neq\tilde{\boldsymbol{w}}_{t_{\ell}}\right)=O(n^{-3})\,.$ (48)
We have from [BVH16], for every $x\geq 0$:
$\mathbb{P}\left(\left\|\tilde{\boldsymbol{w}}_{t_{\ell}}\right\|_{\text{op}}\geq 4\sigma+x\right)\leq n\exp\left(-\dfrac{cx^{2}}{\sigma_{\star}^{2}}\right)\,,$ (49)
for some absolute constant $c>0$, where
$\sigma^{2}$ $:=\max_{i\leq n}\sum_{j=1}^{n}\mathbb{E}\left[(\tilde{\boldsymbol{w}}_{t_{\ell}})_{ij}^{2}\right]\leq\sum_{j=1}^{n}\mathbb{E}\left[\eta_{s}\left(\dfrac{\boldsymbol{w}_{t_{\ell}}}{\sqrt{t_{\ell}}}\right)_{ij}^{2}\right]\,,$ (50)
$\sigma_{\star}$ $:=\max_{i,j\leq n}|(\tilde{\boldsymbol{w}}_{t_{\ell}})_{ij}|\leq 8t_{\ell}^{1/4}\sqrt{\log n}\,.$ (51)
It can be seen from an immediate Gaussian calculation that, for $i\neq j$ and $Z\sim{\mathsf{N}}(0,1)$:
$\mathbb{E}\left[\eta_{s}\left(\dfrac{\boldsymbol{w}_{t_{\ell}}}{\sqrt{t_{\ell}}}\right)_{ij}^{2}\right]=$ $\int_{0}^{\infty}4z\cdot\mathbb{P}\left(\sqrt{\frac{1+\varepsilon}{2}}Z\geq z+s\right){\rm d}z$
$\overset{(a)}{\leq}$ $\int_{0}^{\infty}4z\cdot\dfrac{1}{z+s}\cdot\exp\left(-\dfrac{(z+s)^{2}}{1+\varepsilon}\right){\rm d}z$
$\overset{(b)}{\ll}$ $\dfrac{1}{s}\exp\left(-\dfrac{s^{2}}{1+\varepsilon}\right)\leq\exp\left(-\dfrac{s^{2}}{1+\varepsilon}\right)$
Here in $(a)$ we employ the Mill’s ratio bound, and $(b)$ follows from $z+s\geq s$ and $s\to\infty$.
Proceeding analogously for the diagonal entries of $\eta_{s}(\boldsymbol{w}_{t_{\ell}}/\sqrt{t_{\ell}})$, we obtain that $\sigma\ll\sqrt{n}\exp(-s^{2}/(2(1+\varepsilon)))=k$ by definition of $s$.
We set $x=k/3+\sqrt{t_{\ell}}/(3s)$. Since $x\gg\sigma$, we have $4\sigma+x\leq(3/2)x$ if $n,k$ are sufficiently large. Using Eq. (49) we obtain that, for some universal constants $c,c^{\prime},c^{\prime\prime}>0$:
$\mathbb{P}\left(\|\tilde{\boldsymbol{w}}_{t_{\ell}}\|_{\text{op}}\geq\dfrac{k}{2}+\dfrac{\sqrt{t_{\ell}}}{2s}\right)\leq n\exp\left(-\dfrac{c\left(\dfrac{k}{3}+\dfrac{\sqrt{t_{\ell}}}{3s}\right)^{2}}{64t_{\ell}^{1/2}\log n}\right)\overset{(a)}{\leq}n\exp\left(-\dfrac{c^{\prime}k}{s\log n}-\dfrac{c^{\prime\prime}t_{\ell}^{1/2}}{s^{2}\log n}\right)\,.$
In step $(a)$, we simply expand the squared term in the numerator and drop the quadratic term in $k$. Now, taking a union bound over $\ell\geq 1$, we get that (similar to Eq. (45))
$\mathbb{P}\left(\exists\ell\geq 1:\|\tilde{\boldsymbol{w}}_{t_{\ell}}\|_{\text{op}}\geq\dfrac{k}{2}+\dfrac{\sqrt{t_{\ell}}}{2s}\right)\leq$ $\;n\exp\left(-\dfrac{c^{\prime}k}{s\log n}\right)\sum_{\ell=1}^{\infty}\exp\left(-\dfrac{c^{\prime\prime}t_{\ell}^{1/2}}{s^{2}\log n}\right)$
$\leq$ $\;n\exp\left(-\dfrac{c^{\prime}k}{s\log n}\right)\left(O(1)+\int_{0}^{\infty}\exp\left(-\dfrac{c^{\prime\prime}x}{s^{2}n^{3}\log n}\right){\rm d}x\right)$
$=$ $\;O\left(\exp\left(-\dfrac{c^{\prime}k}{s\log n}\right)\cdot s^{2}n^{4}\log n\right)$
$=$ $\;o_{n}(1)\,,$
where the last estimate holds if $k\geq C(\log n)^{5/2}$ for some sufficiently large $C>0$. In conclusion, using the last display and Eq. (48) we have shown that
$\mathbb{P}\left(\exists\ell\geq 1:\left\|\eta_{s}\left(\dfrac{\boldsymbol{w}_{t_{\ell}}}{\sqrt{t_{\ell}}}\right)\right\|_{\text{op}}\geq\dfrac{k}{2}+\dfrac{\sqrt{t_{\ell}}}{2s}\right)=o_{n}(1)\,.$ (52)
Now we control the in-between fluctuations. Noting that $\eta_{s}(\cdot)$ is a $1$-Lipschitz function, we have the following crude bound:
$\max_{t_{\ell}\leq t\leq t_{\ell+1}}\left\|\eta_{s}\left(\dfrac{\boldsymbol{w}_{t}}{\sqrt{t}}\right)-\eta_{s}\left(\dfrac{\boldsymbol{w}_{t_{\ell}}}{\sqrt{t_{\ell}}}\right)\right\|_{\text{op}}\leq$ $\max_{t_{\ell}\leq t\leq t_{\ell+1}}\left\|\dfrac{\boldsymbol{w}_{t}}{\sqrt{t}}-\dfrac{\boldsymbol{w}_{t_{\ell}}}{\sqrt{t_{\ell}}}\right\|_{F}$
$\leq$ $\dfrac{\max_{t_{\ell}\leq t\leq t_{\ell+1}}\|\boldsymbol{w}_{t}-\boldsymbol{w}_{t_{\ell}}\|_{F}}{\sqrt{t_{\ell}}}+\|\boldsymbol{w}_{t_{\ell}}\|_{F}\left(\dfrac{1}{\sqrt{t_{\ell}}}-\dfrac{1}{\sqrt{t_{\ell+1}}}\right)\,.$
From Lemma A.3, we obtain that
$\mathbb{P}\left(\max_{t_{\ell}\leq t\leq t_{\ell+1}}\left\|\eta_{s}(\boldsymbol{w}_{t}/\sqrt{t})-\eta_{s}(\boldsymbol{w}_{t_{\ell}}/\sqrt{t_{\ell}})\right\|_{\text{op}}\geq 4n\cdot\sqrt{t_{\ell+1}-t_{\ell}}+4n\cdot\sqrt{t_{\ell}}\cdot\left(1-\sqrt{\dfrac{t_{\ell}}{t_{\ell+1}}}\right)\right)\leq 4\,e^{-n^{2}t_{\ell}/4}\,.$
By definition of $t_{\ell}$, simple algebra reveals that (we also use the fact that $n^{-1/2}\ll s^{-1}$):
$4n\cdot\sqrt{t_{\ell+1}-t_{\ell}}+4n\cdot\sqrt{t_{\ell}}\cdot\left(1-\sqrt{\dfrac{t_{\ell}}{t_{\ell+1}}}\right)\leq\\ \dfrac{\sqrt{t_{\ell}}}{2s}\,.$
By union bound over $\ell\geq 1$,
$\mathbb{P}\left(\exists\ell\geq 1:\max_{t_{\ell}\leq t\leq t_{\ell+1}}\left\|\eta_{s}(\boldsymbol{w}_{t}/\sqrt{t})-\eta_{s}(\boldsymbol{w}_{t_{\ell}}/\sqrt{t_{\ell}})\right\|_{\text{op}}\geq\dfrac{k}{2}+\dfrac{\sqrt{t_{\ell}}}{2s}\right)$
$\leq$ $4\sum_{\ell=1}^{\infty}\exp\left(-\dfrac{n^{2}}{8}-\dfrac{t_{\ell}}{8}\right)$
$=$ $4\exp(-n^{2}/8)\sum_{\ell=1}^{\infty}\exp\left(-\dfrac{1}{8}\left(1+\dfrac{\ell-1}{n^{3}}\right)^{2}\right)$
[…]
Using this estimate together with Eq. (52), we conclude that with high probability the following holds simultaneously for all $t\geq 1$. Letting $\ell$ be largest such that $t_{\ell}\leq t$:
$\left\|\eta_{s}(\boldsymbol{w}_{t}/\sqrt{t})\right\|_{\text{op}}\leq\left\|\eta_{s}(\boldsymbol{w}_{t_{\ell}}/\sqrt{t_{\ell}})\right\|_{\text{op}}+\left\|\eta_{s}(\boldsymbol{w}_{t}/\sqrt{t})-\eta_{s}(\boldsymbol{w}_{t_{\ell}}/\sqrt{t_{\ell}})\right\|_{\text{op}}\leq k+\dfrac{\sqrt{t_{\ell}}}{s}\leq k+\dfrac{\sqrt{t}}{s}\,,$
and this finishes the proof. ∎
Remark A.1.
We remark that Assumption (C2) in the proof above is technically not needed, meaning that the additional noise stream $\boldsymbol{g}_{t}$ can in fact be discarded entirely: an appropriate thresholding of $\boldsymbol{v}_{t}$, the top eigenvector of $\eta_{s}(\boldsymbol{A}_{t,+})$, as in Algorithm 1, will also suffice to satisfy all conditions of Theorem 3, although $\boldsymbol{x}$ will not be recovered exactly; some $o(k)$ positions outside the support of $\boldsymbol{x}$ will also be chosen, at most. The reason for this is that the alignment $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|=1-o_{n}(1)$ already, from a close inspection of our proof of Proposition A.1. Regarding the proof of Proposition A.4 above, one can easily realize that even if $\varepsilon=0$, it will go through without any modification. We choose to keep our formulation of Algorithm 2 as faithful to the original work of [CLR17] as possible to discuss a variety of approaches, and leave this to the interested reader.
Appendix B Proof of Proposition 2.1
We take the first row of $\hat{\boldsymbol{z}}_{1}$, and let $A=\{z_{11},\cdots,z_{1n}\}$. Let $r_{j}=\text{rank}(z_{1j})$ denote the rank of $z_{1j}$ with respect to the elements of $A$. Then since $z_{1j}\sim{\mathsf{N}}(0,1)$ across $j$, the collection of the first $k$ ranks $A_{k}=\{r_{1},\cdots,r_{k}\}$ constitutes a sample without replacement from $[n]$. Construct $\boldsymbol{v}$ a binary vector such that $v_{i}=1$ if and only if $i\in A_{k}$, and let $\boldsymbol{u}$ be a randomized-sign vector version of $(1/\sqrt{k})\boldsymbol{v}$. Let
$\hat{\boldsymbol{m}}(\boldsymbol{y},t;\boldsymbol{g}_{1})=\boldsymbol{u}\boldsymbol{u}^{{\sf T}}=\boldsymbol{x}^{\prime}$ (53)
then it is clear that $\hat{\boldsymbol{m}}(\boldsymbol{y},t;\hat{\boldsymbol{z}}_{1})\sim\boldsymbol{x}$ and is independent of $\boldsymbol{x}$ (as it is a function of only $\hat{\boldsymbol{z}}_{1}$). The identity from $(i)$ follows accordingly. To see that this error is clearly sub-optimal compared to polynomial time algorithms, observe that $\hat{\boldsymbol{m}}=\boldsymbol{0}$ is a polynomial time drift, which achieves error $1$ at every $t$.
Point $(ii)$ also follows immediately. Indeed, for every $\ell\geq 0$,
$W_{1}(\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{\ell\Delta},\ell\Delta),\boldsymbol{x})=W_{1}(\boldsymbol{x}^{\prime},\boldsymbol{x})=0$
Lastly, regarding point $(iii)$, note that since $\boldsymbol{x}^{\prime}$ is not dependent on $t$, we have, for every $\ell\geq 1$,
$\hat{\boldsymbol{y}}_{\ell\Delta}=\hat{\boldsymbol{y}}_{(\ell-1)\Delta}+\Delta\boldsymbol{x}^{\prime}+\sqrt{\Delta}\hat{\boldsymbol{z}}_{\ell\Delta}$
Simple induction gives
$\hat{\boldsymbol{y}}_{\ell\Delta}=(\ell\Delta)\boldsymbol{x}^{\prime}+\sqrt{\Delta}\sum_{j=1}^{\ell}\hat{\boldsymbol{z}}_{j\Delta}$
We take the coupling of $(\hat{\boldsymbol{y}}_{\ell\Delta}/(\ell\Delta),\boldsymbol{x})$ such that $\boldsymbol{x}^{\prime}=\boldsymbol{x}$. Then by definition of the Wasserstein-$2$ metric,
$W_{2}(\hat{\boldsymbol{y}}_{\ell\Delta}/(\ell\Delta),\boldsymbol{x})^{2}\leq\mathbb{E}\left[\left\|\dfrac{\sqrt{\Delta}\sum_{j=1}^{\ell}\hat{\boldsymbol{z}}_{j\Delta}}{\ell\Delta}\right\|^{2}\right]=\dfrac{n^{2}}{\ell\Delta}$
It is clear that as $\ell\to\infty$, this quantity converges to 00.
Appendix C Proof of Proposition A.1
We conduct our analysis conditional on $\boldsymbol{u}$ (c.f. Equation (1.2)), and $S$ be the support of $\boldsymbol{u}$. Let $\boldsymbol{A}_{0}=\sqrt{t}\boldsymbol{u}\boldsymbol{u}^{{\sf T}}$ and notice that
$\boldsymbol{A}_{t,+}=\boldsymbol{A}_{0}+\sigma_{+}\boldsymbol{Z}\,,\;\;\;\;\;\boldsymbol{A}_{t,-}=\boldsymbol{A}_{0}+\sigma_{-}\boldsymbol{W}\,,$ (54)
where $\sigma^{2}_{+}:=(1+\varepsilon)/2$, $\sigma^{2}_{-}:=(1+\varepsilon)/(2\varepsilon)$, and $\boldsymbol{Z},\boldsymbol{W}\sim{\sf GOE}(n)$ are independent random matrices.
We have
$\eta_{s}(\boldsymbol{A}_{t,+})=\boldsymbol{A}_{0}+\eta_{s}(\sigma_{+}\boldsymbol{Z})+\mathbb{E}[\boldsymbol{B}]+\big{(}\boldsymbol{B}-\mathbb{E}[\boldsymbol{B}]\big{)}\,,$ (55)
where
$B_{ij}=\eta_{s}\Big{(}\sqrt{t}\cdot u_{i}u_{j}+\sigma_{+}Z_{ij}\Big{)}-\sqrt{t}\cdot u_{i}u_{j}-\eta_{s}(\sigma_{+}Z_{ij})\,.$ (56)
Our first order of business is to analyze $\mathbb{E}[\boldsymbol{B}]$. If $i\notin S$ or $j\notin S$, we have $\mathbb{E}[B_{ij}]=0$. On the other hand, if $i,j\in S$, then
  • (i)
    Case 1: $u_{i}u_{j}=1/k$
    In this case, we have $\mathbb{E}[B_{ij}]=-b_{0}+b_{1}\boldsymbol{1}_{i=j}$ where (below $G\sim{\mathsf{N}}(0,1)$)
    $b_{0}:=-\mathbb{E}\Big{\{}\eta_{s}\Big{(}\frac{\sqrt{t}}{k}+\sigma_{+}G\Big{)}-\frac{\sqrt{t}}{k}\Big{\}}\,,\;\;\;b_{1}:=\mathbb{E}\Big{\{}\eta_{s}\Big{(}\frac{\sqrt{t}}{k}+\sqrt{2}\sigma_{+}G\Big{)}-\eta_{s}\Big{(}\frac{\sqrt{t}}{k}+\sigma_{+}G\Big{)}\Big{\}}\,.$ (57)
    Recalling that $\sigma_{+}$ is bounded and bounded away from 00 (without loss of generality we can assume $\varepsilon<1/2$) and $s,\sqrt{t}/k-s$ grows with $n,k$, so that $\eta_{s}(\sqrt{t}/k+Z_{ij})=\sqrt{t}/k+Z_{ij}-s$ with high probability; hence $|B_{ij}+s|=o_{P}(s)$ (as $Z_{ij}=o(s)$ with high probability). Noting that $|B_{ij}|\leq 2s$, we get $b_{0}=s(1+o(1))$ and $b_{1}=o(s)$ (distribution on diagonal is different).
  • (ii)
    Case 2: $u_{i}u_{j}=-1/k$
    By a similar reasoning, we have $\mathbb{E}[B_{ij}]=s(1+o(1))$.
We can thus rewrite
$\eta_{s}(\boldsymbol{A}_{t,+})=(\sqrt{t}-kb_{0})\cdot\boldsymbol{u}\boldsymbol{u}^{{\sf T}}+b_{1}\cdot{\boldsymbol{P}}_{S}+\eta_{s}(\sigma_{+}\boldsymbol{Z})+\big{(}\boldsymbol{B}-\mathbb{E}[\boldsymbol{B}]\big{)}\,,$ (58)
where $(\boldsymbol{P}_{S})_{ij}=1$ if $i=j\in S$ and $=0$ otherwise.
Next, we analyze the operator norm of $\eta_{s}(\sigma_{+}\boldsymbol{Z})$. Let $\tilde{\boldsymbol{Z}}=(\tilde{Z}_{ij})_{i,j\leq n}$ be defined as
[…]
for some constant $C>0$ that we choose such that $\max_{ij\leq n}|Z_{ij}|\leq C\sqrt{\log n}$ with probability at least $1-n^{-D}$ ($D$ is defined in the statement of Proposition A.1). We have $\tilde{\boldsymbol{Z}}=\eta_{s}(\boldsymbol{Z})$. By [BVH16], there exists an absolute constant $c>0$ such that for every $u>0$:
$\mathbb{P}\left(\|\tilde{\boldsymbol{Z}}\|_{{\mbox{\rm\tiny op}}}\geq 4\sigma+u\right)\leq n\exp\left(-\frac{cu^{2}}{L^{2}}\right)\,,$ (60)
where
$\sigma^{2}$ $=\max_{i\leq n}\sum_{j=1}^{n}\mathbb{E}[\eta_{s}(Z_{ij})^{2}]\,,$ (61)
$L$ $=\max_{i,j\leq n}\|\tilde{Z}_{ij}\|_{\infty}\leq C\sqrt{\log n}\,.$ (62)
An immediate Gaussian calculation yields, for $i\neq j$:
$\mathbb{E}[\eta_{s}(\sigma_{+}Z_{ij})^{2}]=$ $\int_{0}^{\infty}4z\cdot\mathbb{P}(\sigma_{+}Z_{ij}\geq z+s){\rm d}z\leq C_{1}e^{-s^{2}/(1+\varepsilon)}\,.$ (63)
for some constant $C_{1}>0$.
Proceeding analogously for $\eta_{s}(\sigma_{+}Z_{ii})$ and substituting in Eq. (60), we get $\sigma^{2}\leq 2C_{1}n\exp\{-s^{2}/(1+\varepsilon)\}$. Applying Eq. (60) there exists an absolute constant $C>0$ such that, with probability at least $1-2n^{-D}$,
$\|\eta_{s}(\boldsymbol{Z})\|_{{\mbox{\rm\tiny op}}}$ $\leq C\Big{(}\sqrt{n}\exp\{-s^{2}/2(1+\varepsilon)\}\vee\sqrt{\log n}\Big{)}\,$ (64)
$\leq C\Big{(}k\vee\sqrt{\log n}\Big{)}\leq Ck\,.$ (65)
Lastly, consider $\boldsymbol{B}-\mathbb{E}[\boldsymbol{B}]$. By Eq. (56) we know that the entries of this matrix are independent with mean 00 and bounded by $2s$, hence subgaussian. Further only a $k\times k$ submatrix is nonzero, so that
$\|\boldsymbol{B}-\mathbb{E}[\boldsymbol{B}]\|_{{\mbox{\rm\tiny op}}}\leq C_{1}\sqrt{k}s\,,$ (66)
with high probability (for instance, the operator norm tail bound above can be applied once more, which gives an error probability of at most $C\exp(-ck)$ for some absolute constant $C,c>0$).
Summarizing, we proved that
$\eta_{s}(\boldsymbol{A}_{t,+})$ $=(\sqrt{t}-kb_{0})\cdot\boldsymbol{u}\boldsymbol{u}^{{\sf T}}+\boldsymbol{\Delta}\,,$ (67)
$\|\boldsymbol{\Delta}\|_{{\mbox{\rm\tiny op}}}$ $\leq C(k+\sqrt{k}s)\leq C^{\prime}k\,,$ (68)
where in the last step we used $k\gg(\log n)^{2}$
Recall that $\boldsymbol{v}_{t}$ denotes the top eigenvector of $\eta_{s}(\boldsymbol{A}_{t,+})$. By Davis-Kahan,
$\min_{a\in\{+1,-1\}}\left\|\boldsymbol{v}_{t}-a\boldsymbol{u}\right\|$ $\leq\dfrac{Ck}{\sqrt{t}-kb_{0}}$ (69)
$\stackrel{{\scriptstyle(a)}}{{\leq}}\dfrac{Ck}{\sqrt{t}-(1+\varepsilon)ks}$ (70)
$\stackrel{{\scriptstyle(b)}}{{\leq}}\varepsilon\,,$ (71)
where in $(a)$ we used the fact that $b_{0}=s+o(s)$ and in $(b)$ the fact that $t\geq(1+\delta)k^{2}\log(n/k^{2})$, whereby we can assume $\delta\geq C\varepsilon$ for $C$ a sufficiently large absolute constant. Recalling the definition of the score $\hat{\boldsymbol{v}}_{t}$:
$\hat{\boldsymbol{v}}_{t}=\boldsymbol{A}_{t,-}\boldsymbol{v}_{t}=\sqrt{t}\boldsymbol{u}\langle\boldsymbol{u},\boldsymbol{v}_{t}\rangle+\sigma_{-}\boldsymbol{W}\boldsymbol{v}_{t}$
where we know that $\boldsymbol{G}=\boldsymbol{W}\boldsymbol{v}_{t}\sim{\mathsf{N}}(0,\boldsymbol{I}_{n})$ by independence of $\boldsymbol{W}$ and $\boldsymbol{v}_{t}$. Assuming to be definite that the sign of the eigenvector is chosen so that the last bound holds with $a=+1$, we get that $\langle\boldsymbol{u},\boldsymbol{v}_{t}\rangle\geq 1-\varepsilon^{2}$. We get that for every $i\in S$:
$u_{i}>0$ $\Rightarrow\hat{v}_{t,j}\geq(1-\varepsilon^{2})\dfrac{\sqrt{t}}{k}-\sigma_{-}|G_{j}|\geq(1-\varepsilon^{2})\dfrac{\sqrt{t}}{k}-\dfrac{4}{\varepsilon}\sqrt{\log n}$ (72)
$u_{i}<0$ $\Rightarrow\hat{v}_{t,j}\leq-(1-\varepsilon^{2})\dfrac{\sqrt{t}}{k}+\sigma_{-}|G_{j}|\leq-(1-\varepsilon^{2})\dfrac{\sqrt{t}}{k}+\dfrac{4}{\varepsilon}\sqrt{\log n}$ (73)
where we use a union bound to get $|G_{j}|\leq 4\sqrt{\log n}$ for all $j\leq n$. Similarly, for all $i\notin S$,
$|\hat{v}_{t,j}|\leq\sigma_{-}|G_{j}|\leq\dfrac{4}{\varepsilon}\sqrt{\log n}$
These calculations reveal that: (i) the entries with the largest magnitudes are the elements of $S$, and (ii) if $u_{i}$ and $\hat{v}_{t,i}$ share the same sign for all $i\in S$. On this event, $\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|=0$.
Lastly, we claim that the top eigenvalue of $\eta_{s}(\boldsymbol{A}_{t,+})$ is larger than $k+\sqrt{t}/s$. From triangle inequality applied to Eq. (58), we have
$\lambda_{1}(\eta_{s}(\boldsymbol{A}_{t,+}))$ $\geq(\sqrt{t}-kb_{0})-b_{1}-\|\eta_{s}(\sigma_{+}\boldsymbol{Z})\|_{{\mbox{\rm\tiny op}}}-\|\boldsymbol{B}-\mathbb{E}[\boldsymbol{B}]\|_{{\mbox{\rm\tiny op}}}$ (74)
$\stackrel{{\scriptstyle(a)}}{{\geq}}(\sqrt{t}-(1+\varepsilon)ks)-Ck-C\sqrt{k}s$ (75)
$\stackrel{{\scriptstyle(b)}}{{\geq}}\sqrt{t}-(1+C_{0}\varepsilon)ks$ (76)
$\stackrel{{\scriptstyle(c)}}{{\geq}}\varepsilon ks\,.$ (77)
Here $(a)$ follows from Eqs. (65) and (66), $(b)$ because $k\gg 1$ and $s\gg 1$ and $(c)$ follows by taking $\delta\geq C\varepsilon$ for $C$ a sufficiently large absolute constant. The claim follows because $ks\gg k$ and $ks\gg\sqrt{t}$.
Appendix D Proof of Lemma A.3
Proof.
[…]
$\mathbb{P}\left(\max_{0\leq t\leq t_{\ell+1}-t_{\ell}}\|\boldsymbol{w}_{t}\|_{F}-\mathbb{E}\left[\max_{0\leq t\leq t_{\ell+1}-t_{\ell}}\|\boldsymbol{w}_{t}\|_{F}\right]\geq x\right)\leq 2\exp\left(-\dfrac{x^{2}}{4(t_{\ell+1}-t_{\ell})}\right)$
This can be proven, e.g. by discretizing the interval $[0,t_{\ell+1}-t_{\ell}]$ into $r$ equal-length intervals and employing standard Gaussian concentration on vectors (then pushing $r\to\infty$). As the argument is standard, we omit the proof for brevity.
Now we bound $\mathbb{E}\left[\max_{0\leq t\leq t_{\ell+1}-t_{\ell}}\|\boldsymbol{w}_{t}\|_{F}\right]$. We know that $\|\boldsymbol{w}_{t}\|_{F}$ is a non-negative submartingale, so that from Doob’s inequality:
$\mathbb{E}\left[\max_{0\leq t\leq t_{\ell+1}-t_{\ell}}\|\boldsymbol{w}_{t}\|_{F}^{2}\right]\leq 4\mathbb{E}[\|\boldsymbol{w}_{t_{\ell+1}-t_{\ell}}\|_{F}^{2}]\leq 9(t_{\ell+1}-t_{\ell})n^{2}$
so that from Cauchy-Schwarz, $\mathbb{E}\left[\max_{0\leq t\leq t_{\ell+1}-t_{\ell}}\|\boldsymbol{w}_{t}\|_{F}\right]\leq 3\sqrt{t_{\ell+1}-t_{\ell}}n$. Hence as $t_{\ell}\geq 1$,
$\mathbb{P}\left(\max_{t_{\ell}\leq t\leq t_{\ell+1}}\|\boldsymbol{w}_{t}-\boldsymbol{w}_{t_{\ell}}\|_{F}\geq 4\sqrt{(t_{\ell+1}-t_{\ell})t_{\ell}}\cdot n\right)\leq 2\exp\left(-\dfrac{n^{2}t_{\ell}}{4}\right)$
The second tail bound follows immediately (at least, the proof would be analogous to the preceding display). ∎
Appendix E Auxiliary lemmas for Section 5
E.1 Proof of Lemma 5.1
We let $\boldsymbol{B}\sim{\mathsf{N}}(\boldsymbol{0},\boldsymbol{I}_{n^{2}})$ so that $\boldsymbol{W}=(\boldsymbol{B}+\boldsymbol{B}^{{\sf T}})/2$. For $\boldsymbol{v}\in\Omega_{n,k}$ we have $\langle\boldsymbol{v},\boldsymbol{W}\boldsymbol{v}\rangle=\langle\boldsymbol{v},\boldsymbol{B}\boldsymbol{v}\rangle\sim{\mathsf{N}}(0,1)$. We thus have, by Gaussian tail bounds and a triangle inequality:
$\mathbb{P}\left(|\langle\boldsymbol{v},\boldsymbol{W}\boldsymbol{v}\rangle|\geq C\sqrt{\log{\binom{n}{k}}}\right)\leq 2\exp\left(-\dfrac{C^{2}}{2}\log\binom{n}{k}\right)\,.$
Taking the union bound over $\boldsymbol{v}\in\Omega_{n,k}$ gives the desired statement, since the cardinality of this set is $\binom{n}{k}2^{k}$.
E.2 Proof of Lemma 5.3
Using Lemma 5.2, we can take $x=n^{-1/4}$, say, and $\theta=\sqrt{1+\delta}$ for $\delta\geq n^{-c_{0}}$ for some small enough $c_{0}>0$, to get that
$\mathbb{P}\left(\lambda_{1}(\boldsymbol{y})\leq\theta+1/\theta-n^{-1/4}-2/n\right)\leq C\exp(-cn^{1/3})\,,$
for some absolute constants $C,c>0$.
We have the following identity, letting $\boldsymbol{W}\sim{\sf GOE}(n,1/n)$:
$\langle\boldsymbol{u},\boldsymbol{v}_{1}\rangle^{2}=\dfrac{1}{\theta^{2}\langle\boldsymbol{u},(\lambda_{1}(\boldsymbol{y})\boldsymbol{I}-\boldsymbol{W})^{-2}\boldsymbol{u}\rangle}\geq\dfrac{1}{\theta^{2}\cdot\|(\lambda_{1}(\boldsymbol{y})\boldsymbol{I}-\boldsymbol{W})^{-2}\|_{{\mbox{\rm\tiny op}}}}\,.$
By standard Gaussian concentration, we know that, for any $\Delta>0$
$\mathbb{P}\left(\|\boldsymbol{W}\|_{{\mbox{\rm\tiny op}}}\geq 2+\Delta\right)\leq C\exp(-cn\Delta^{2})\,.$
In this inequality, we take
$\Delta=\frac{1}{4}\big{(}\theta+1/\theta-n^{-1/4}-2/n-2\big{)}\,.$
Note that with $\theta=\sqrt{1+\delta}$ and $\delta=o_{n}(1)$, we know that $\theta+1/\theta-2=\Theta(\delta^{2})$, so that $\Delta=\Theta(\delta^{2})$ if $\delta\geq n^{-c_{0}}$ with $c_{0}\leq 1/8$. Hence, by a union bound on the two concentration inequalities,
$\mathbb{P}\left(\lambda_{\min}(\lambda_{1}(\boldsymbol{y})\boldsymbol{I}-\boldsymbol{W})\leq 2\Delta\right)\leq C\exp(-cn^{1/3})$
and on the complement of this event, we know that
$\langle\boldsymbol{v}_{1},\boldsymbol{u}\rangle^{2}\geq\dfrac{4\Delta^{2}}{\theta^{2}}=\Theta(\delta^{4})$
since $\theta=\Omega(1)$, and so $|\langle\boldsymbol{v}_{1},\boldsymbol{u}\rangle|=\Omega(\delta^{2})$.
Appendix F Proof of Proposition 6.1
In our proof, we will use the following elementary facts.
Fact F.1.
For any deterministic unit vector $\boldsymbol{u}$, a unit vector $\boldsymbol{v}$ is uniformly random on the orthogonal subspace to $\boldsymbol{u}$ if and only if $\langle\boldsymbol{v},\boldsymbol{u}\rangle=0$ and $\boldsymbol{v}\stackrel{{\scriptstyle{\rm d}}}{{=}}\boldsymbol{Q}\boldsymbol{v}$ for every orthogonal matrix $\boldsymbol{Q}$ such that $\boldsymbol{Q}\boldsymbol{u}=\boldsymbol{u}$.
Fact F.2.
Let $\boldsymbol{A}$ be a symmetric matrix, and $\boldsymbol{u}$ a unit vector. Denote $\boldsymbol{B}_{\alpha}=\alpha\boldsymbol{u}\boldsymbol{u}^{{\sf T}}+\boldsymbol{A}$, and let $\boldsymbol{v}(\alpha)$ be a top eigenvector of $\boldsymbol{B}_{\alpha}$. Then $f(\alpha)=|\langle\boldsymbol{v}(\alpha),\boldsymbol{u}\rangle|$ is an increasing function of $\alpha>0$.
Let $\boldsymbol{u}\sim\operatorname{Unif}(\Omega_{n,k})$. Recall that $\boldsymbol{A}_{t}=\sqrt{t}\boldsymbol{u}\boldsymbol{u}^{{\sf T}}+\sqrt{t}\boldsymbol{W}$ where $\boldsymbol{W}\sim{\sf GOE}(n,1/2)$.
We conduct our analysis conditional on $\boldsymbol{u}$. Let $\boldsymbol{v}_{t}$ be a top eigenvector of $\boldsymbol{A}_{t}$. For $t=(1+\delta)n/2$, $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|\overset{a.s.}{\to}\sqrt{\delta/(1+\delta)}$, so that with high probability $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|\geq\sqrt{\delta}/(2\sqrt{1+\delta})$. If $t\geq(1+\delta)n/2$, we can use Fact F.2 to obtain the same result. By choosing $\varepsilon$ such that $2\varepsilon<\sqrt{\delta/(1+\delta)}$, we know from standard concentration of the alignment (Lemma 6.6) that $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|\geq 2\varepsilon$ with probability at least $1-\exp(-cn)$ for some $c>0$ possibly dependent on $(\varepsilon,\delta)$.
By rotational invariance of $\boldsymbol{W}$, $\boldsymbol{w}_{t}:=\dfrac{\boldsymbol{v}_{t}-\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle\boldsymbol{u}}{\|\boldsymbol{v}_{t}-\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle\boldsymbol{u}\|}$ is uniformly random on the orthogonal subspace to $\boldsymbol{u}$. hence, there exists $\boldsymbol{g}\sim{\mathsf{N}}(\boldsymbol{0},\boldsymbol{I}_{n})$, such that
$\boldsymbol{w}_{t}\sim\dfrac{(\boldsymbol{I}_{n}-\boldsymbol{u}\boldsymbol{u}^{{\sf T}})\boldsymbol{g}}{\|(\boldsymbol{I}_{n}-\boldsymbol{u}\boldsymbol{u}^{{\sf T}})\boldsymbol{g}\|}\,.$
Since $\|(\boldsymbol{I}_{n}-\boldsymbol{u}\boldsymbol{u}^{{\sf T}})\boldsymbol{g}\|\sim\|\boldsymbol{g}^{\prime}\|$ for some $\boldsymbol{g}^{\prime}\sim{\mathsf{N}}(\boldsymbol{0},\boldsymbol{I}_{n-1})$, we have, for some constant $c>0$,
$\mathbb{P}\left(\|(\boldsymbol{I}_{n}-\boldsymbol{u}\boldsymbol{u}^{{\sf T}})\boldsymbol{g}\|\leq\dfrac{\sqrt{n}}{2}\right)\leq\exp\left(-cn\right)\,.$
Further, for every $1\leq i\leq n$, we know that
$|((\boldsymbol{I}_{n}-\boldsymbol{u}\boldsymbol{u}^{{\sf T}})\boldsymbol{g})_{i}|\leq|g_{i}|+\|\boldsymbol{u}\|_{\infty}\cdot|\langle\boldsymbol{u},\boldsymbol{g}\rangle|\leq|g_{i}|+\dfrac{|\langle\boldsymbol{u},\boldsymbol{g}\rangle|}{\sqrt{k}}\,.$
We next show that, with the claimed probability, only a few entries of $\boldsymbol{w}_{t}$ can have large magnitude. As a result, less than $\ell$ entries of $\boldsymbol{u}$ can be estimated incorrectly (with $\ell=1$ if $k\ll n/\log n$).
Define $\ell=\lceil n\exp(-a_{n}\cdot n/k)\rceil\geq 1$ (with $a_{n}$ a sequence to be chosen later) and $g^{\text{abs}}_{(\ell)}$ as the $\ell$-th largest value among the $|g_{i}|$’s. We have
$\mathbb{P}\left(|\langle\boldsymbol{u},\boldsymbol{g}\rangle|\geq\sqrt{na_{n}}\right)\leq\exp\left(-\dfrac{na_{n}}{2}\right)\,.$
Furthermore, from a union bound, we get that
$\mathbb{P}\left(g_{(\ell)}^{\text{abs}}\geq\dfrac{2\sqrt{na_{n}}}{\sqrt{k}}\right)$ $\leq{\binom{n}{\ell}}\cdot\exp\left(-\dfrac{2n\ell\cdot a_{n}}{k}\right)$
$\leq\left(\dfrac{en}{\ell}\right)^{\ell}\exp\left(-\dfrac{2n\ell\cdot a_{n}}{k}\right)$
$=\exp\left(-\dfrac{2n\ell\cdot a_{n}}{k}+\ell\log\dfrac{n}{\ell}+\ell\right)\,.$
By definition, we know that $\ell\geq\max\left\{n\exp(-a_{n}\cdot n/k),1\right\}$, so that
$\dfrac{2na_{n}}{k}-\log\dfrac{n}{\ell}-1\geq\dfrac{2na_{n}}{k}-\min\left\{\log n,a_{n}\cdot n/k\right\}-1\geq\dfrac{na_{n}}{2k}$
as long as $na_{n}\gg k$. This means that
$\mathbb{P}\left(g_{(\ell)}^{\text{abs}}\geq\dfrac{2\sqrt{na_{n}}}{\sqrt{k}}\right)\leq\exp\left(-\dfrac{na_{n}}{2k}\right)\,.$
Define the set
${\mathcal{A}}_{n}(t):=\Big{\{}i\leq n:\,|w_{ti}|\geq 6\sqrt{\frac{a_{n}}{k}}\Big{\}}\,.$
By the bounds above we have
$\mathbb{P}\left(|{\mathcal{A}}_{n}(t)|\leq\ell-1\right)\geq 1-e^{-cn}-e^{-na_{n}/2k}-e^{-na_{n}/2}\,.$
Suppose that $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|\geq 2\varepsilon$ also holds, and suppose without loss of generality that $\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle\geq 2\varepsilon$. Then, we have (as long as $6\sqrt{a_{n}}\leq(9/10)\varepsilon$)
$i\in S,i\notin{\mathcal{A}}_{n}(t)\,u_{i}>0\Rightarrow v_{ti}\geq\dfrac{\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle}{\sqrt{k}}-\dfrac{6\sqrt{a_{n}}}{\sqrt{k}}>\dfrac{\varepsilon}{\sqrt{k}}\Rightarrow i\in\hat{S},\operatorname{sign}(v_{ti})>0,,$
$i\in S,i\notin{\mathcal{A}}_{n}(t),u_{i}<0\Rightarrow v_{ti}\leq-\dfrac{\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle}{\sqrt{k}}+\dfrac{6\sqrt{a_{n}}}{\sqrt{k}}<-\dfrac{\varepsilon}{\sqrt{k}}\Rightarrow i\in\hat{S},\operatorname{sign}(v_{ti})<0\,.$
Analogously, for $i\notin S,i\notin{\mathcal{A}}_{n}(t)$, we have
$|v_{ti}|\leq\dfrac{6\sqrt{a_{n}}}{\sqrt{k}}<\dfrac{\varepsilon}{\sqrt{k}}\,.$
and we obtain that at most $\ell-1$ positions could be mis-identified.
Next, we show that the termination condition (Line 5, Algorithm 1) does not trigger for each $t\geq n^{2}$ (with high probability). We write
$\boldsymbol{A}_{t}=\dfrac{\boldsymbol{y}_{t}+\boldsymbol{y}_{t}^{{\sf T}}}{2\sqrt{t}}=\sqrt{t}\boldsymbol{u}\boldsymbol{u}^{{\sf T}}+\left(\dfrac{\boldsymbol{B}_{t}+\boldsymbol{B}_{t}^{{\sf T}}}{2\sqrt{t}}\right)$
From Weyl’s inequality:
$\lambda_{1}(\boldsymbol{A}_{t})\geq\sqrt{t}-\left\|\dfrac{\boldsymbol{B}_{t}+\boldsymbol{B}_{t}^{{\sf T}}}{2\sqrt{t}}\right\|_{\text{op}}$
From standard operator norm results for GOE matrices (as $(\boldsymbol{B}_{t}+\boldsymbol{B}_{t}^{{\sf T}})/\sqrt{2t}\sim{\sf GOE}(n)$), we know that $\|(\boldsymbol{B}_{t}+\boldsymbol{B}_{t}^{{\sf T}})/(2\sqrt{t})\|_{\text{op}}\leq 2\sqrt{n}$ with probability at least $1-\exp(-cn)$, for some $c>0$. Hence $\lambda_{1}(\boldsymbol{A}_{t})\geq\sqrt{t}-2\sqrt{n}>\sqrt{t}/2$ as $t\geq n^{2}\gg n$.
We obtain that
$\mathbb{E}\left[\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\right]=O\left(\dfrac{\ell-1}{k}+\exp\left(-\dfrac{na_{n}}{k}\right)\right)=O\left(\exp(-n\varepsilon^{2}/64k)\right)$
where we picked $a_{n}>\varepsilon^{2}/64$ satisfying the bounds outlined above, namely $(i)$ $6\sqrt{a_{n}}\leq 0.9\varepsilon$, and $(ii)$ $na_{n}\gg k$. Notice that $na_{n}/k\gg\log(na_{n}/k)$ if $na_{n}\gg k$, and $\ell-1\leq n\cdot\exp(-a_{n}\cdot n/k)$.
Appendix G Proof of Lemma 6.3
By Gaussian concentration, we have
$\mathbb{P}\left(\max_{0\leq t\leq t_{\ell+1}-t_{\ell}}\|\boldsymbol{W}_{t}\|_{\text{op}}-\mathbb{E}\left[\max_{0\leq t\leq t_{\ell+1}-t_{\ell}}\|\boldsymbol{W}_{t}\|_{\text{op}}\right]\geq x\right)\leq 2\exp\left(-\dfrac{x^{2}}{2(t_{\ell+1}-t_{\ell})}\right)$
This can be proven, e.g. by discretizing the interval $[0,t_{\ell+1}-t_{\ell}]$ into $r$ equal-length intervals and employing Gaussian concentration on vectors (then pushing $r\to\infty$). As the argument is standard, we omit the proof for brevity.
To evaluate $\mathbb{E}[\max_{0\leq t\leq t_{\ell+1}-t_{\ell}}\|\boldsymbol{W}_{t}\|_{\text{op}}]$, we recognize that $\|\boldsymbol{W}_{t}\|_{\text{op}}$ is a submartingale, so that from Doob’s inequality:
$\mathbb{E}\left[\max_{0\leq t\leq t_{\ell+1}-t_{\ell}}\|\boldsymbol{W}_{t}\|_{\text{op}}^{2}\right]\leq 4\mathbb{E}\left[\|\boldsymbol{W}_{t_{\ell+1}-t_{\ell}}\|_{\text{op}}^{2}\right]$
Once again from Gaussian concentration,
$\mathbb{P}\left(|\|\boldsymbol{W}_{1}\|_{\text{op}}-\mathbb{E}[\|\boldsymbol{W}_{1}\|_{\text{op}}]|\geq x\right)\leq 2\exp\left(\dfrac{-x^{2}}{2}\right)$
so that $\mathbb{P}\left(|\|\boldsymbol{W}_{t_{\ell+1}-t_{\ell}}\|_{\text{op}}-\mathbb{E}[\|\boldsymbol{W}_{t_{\ell+1}-t_{\ell}}\|_{\text{op}}]|\geq x\right)\leq 2\exp\left(-x^{2}/(2(t_{\ell+1}-t_{\ell}))\right)$. Hence $\|\boldsymbol{W}_{t_{\ell+1}-t_{\ell}}\|_{\text{op}}$ is $(t_{\ell+1}-t_{\ell})$-subgaussian, implying that $\text{Var}(\|\boldsymbol{W}_{t_{\ell+1}-t_{\ell}}\|_{\text{op}})\leq 6(t_{\ell+1}-t_{\ell})$. As $\mathbb{E}[\|\boldsymbol{W}_{t_{\ell+1}-t_{\ell}}\|_{\text{op}}]^{2}\sim 4(t_{\ell+1}-t_{\ell})n$ (one can obtain this from the Bai-Yin Theorem along with sub-gaussianity, for instance), we get that $\mathbb{E}\left[\|\boldsymbol{W}_{t_{\ell+1}-t_{\ell}}\|^{2}\right]\leq 16(t_{\ell+1}-t_{\ell})n$ eventually as $n$ gets large.
From Cauchy-Schwarz inequality, we get that
$\mathbb{E}\left[\max_{0\leq t\leq t_{\ell+1}-t_{\ell}}\|\boldsymbol{W}_{t}\|_{\text{op}}\right]\leq 8\sqrt{t_{\ell+1}-t_{\ell}}\sqrt{n}$
We conclude that
$\mathbb{P}\left(\max_{0\leq t\leq t_{\ell+1}-t_{\ell}}\|\boldsymbol{W}_{t}\|_{\text{op}}\geq 16\sqrt{(t_{\ell+1}-t_{\ell})n}\right)\leq 2\exp\left(-32n\right)$
Appendix H Proof of Lemma 6.4
First, by orthogonal invariance of $\boldsymbol{W}_{t}$, we know that $\boldsymbol{v}_{t}$ is uniformly random over the unit sphere $\mathbb{S}^{n-1}$. We can write, using $\boldsymbol{g}\sim{\mathsf{N}}(0,\boldsymbol{I}_{n})$, the following representation
$\boldsymbol{v}_{t}\sim\dfrac{\boldsymbol{g}}{\|\boldsymbol{g}\|}$
As in the statement of the Lemma, we define the following set, for $\boldsymbol{v}\in\mathbb{R}^{n}$ and $C>0$:
$A(\boldsymbol{v};C)=\left\{i:1\leq i\leq n,|v_{i}|\geq\dfrac{C\sqrt{\log(n/k)}}{\sqrt{n}}\right\}$
As with the proof of Proposition 6.1, we first deal with the denominator $\|\boldsymbol{g}\|$: indeed, sub-exponential concentration gives us
$\mathbb{P}\left(\sum_{j=1}^{n}g_{j}^{2}\leq\dfrac{n}{2}\right)$ $\leq 2\exp(-n/8)$ (78)
This leads us to define another set
$B(\boldsymbol{g};C)=\left\{i:1\leq i\leq n,|g_{i}|\geq C\sqrt{\log(n/k)}\right\}$
Let $p_{n}=\mathbb{P}(|g_{1}|\geq C\sqrt{\log(n/k)})$, then we have $|B(\boldsymbol{g};C)|\sim\text{Bin}(n,p_{n})$. From Gaussian tail bounds, we know that $p_{n}\leq(n/k)^{-C^{2}/2}$. We now use a Chernoff bound of the following form: for every $x\geq 4\mathbb{E}[X]$, where $X\sim\text{Bin}(n,p)$, then
$\mathbb{P}\left(X\geq x\right)\leq\exp\left(-x/3\right)$
It is clear that $np_{n}\ll k^{2}/n\leq\max\{k^{2}/n,\sqrt{k}\}$ when $C>2$, so that we have
$\mathbb{P}\left(|B(\boldsymbol{g};C)|\geq\max\{\sqrt{k},k^{2}/n\}\right)\leq\exp\left(-\dfrac{1}{3}\max\{\sqrt{k},k^{2}/n\}\right)\leq\exp\left(-\dfrac{1}{3}n^{1/4}\right)$
Therefore, with each fixed $t$, by union bound with probability at least $1-O(\exp(-\sqrt{n}))$, we have, for a possibly different $C>0$, $|A(\boldsymbol{v}_{t};C)|\leq\max\{\sqrt{k},k^{2}/n\}$. Our proof ends here, as $\max\{\sqrt{k},k^{2}/n\}\ll k/2$ for $\sqrt{n}\ll k\ll n$.
Appendix I Proof of Lemma 6.5
We know that
$\boldsymbol{v}_{t}^{{\sf T}}\boldsymbol{W}_{t_{\ell}}\boldsymbol{v}_{t}=$ $\boldsymbol{v}_{t}^{{\sf T}}\boldsymbol{W}_{t}\boldsymbol{v}_{t}-\boldsymbol{v}_{t}^{{\sf T}}(\boldsymbol{W}_{t}-\boldsymbol{W}_{t_{\ell}})\boldsymbol{v}_{t}$
$=$ $\lambda_{1}(\boldsymbol{W}_{t})-\boldsymbol{v}_{t}^{{\sf T}}(\boldsymbol{W}_{t}-\boldsymbol{W}_{t_{\ell}})\boldsymbol{v}_{t}$
$=$ $\lambda_{1}(\boldsymbol{W}_{t_{\ell}})-\boldsymbol{v}_{t}^{{\sf T}}(\boldsymbol{W}_{t}-\boldsymbol{W}_{t_{\ell}})\boldsymbol{v}_{t}+(\lambda_{1}(\boldsymbol{W}_{t})-\lambda_{1}(\boldsymbol{W}_{t_{\ell}}))$
from which we obtain from Weyl’s inequality that
$\sup_{t_{\ell}\leq t\leq t_{\ell+1}}\left|\boldsymbol{v}_{t}^{{\sf T}}\boldsymbol{W}_{t_{i}}\boldsymbol{v}_{t}-\lambda_{1}(\boldsymbol{W}_{t_{\ell}})\right|\leq 2\|\boldsymbol{W}_{t}-\boldsymbol{W}_{t_{\ell}}\|_{\text{op}}\leq 16\sqrt{(t_{\ell+1}-t_{\ell})n}$
with probability at least $1-2\exp(-32n)$.
Appendix J Proof of Lemma 6.6
By Weyl’s inequality, $\boldsymbol{W}\mapsto\lambda_{1}(\boldsymbol{Y})$ (with $\boldsymbol{Y}=\theta\boldsymbol{v}\boldsymbol{v}^{{\sf T}}+\boldsymbol{W}$) is a 1-Lipschitz function and therefore, by Borell inequality (and [BBAP05]), letting $\lambda_{*}(\theta):=\theta+1/\theta$, for any $\varepsilon>0$,
$\mathbb{P}\big{(}|\lambda_{1}(\boldsymbol{Y})-\lambda_{*}(\theta)|\geq\varepsilon\big{)}\leq 2e^{-n\varepsilon^{2}/4}\,.$ (79)
To prove concentration of $\langle\boldsymbol{v}_{1}(\boldsymbol{Y}),\boldsymbol{v}\rangle^{2}$, note that simple linear algebra yields
$\frac{1}{\langle\boldsymbol{v}_{1}(\boldsymbol{Y}),\boldsymbol{v}\rangle^{2}}=\langle\boldsymbol{v},(\lambda_{1}(\boldsymbol{Y})\boldsymbol{I}-\boldsymbol{W})^{-2}\boldsymbol{v}\rangle=:F(\boldsymbol{W})\,.$ (80)
It is therefore sufficient to prove that $F(\boldsymbol{W})$ concentrates around a value that is bounded away from 00. Fix $\varepsilon_{0}>0$ such that $2+3\varepsilon_{0}<\lambda_{*}(\theta)$ and define the event
:=$\mathcal{E}:=\big{\{}\boldsymbol{W}:\,\|\boldsymbol{W}\|_{{\mbox{\rm\tiny op}}}\leq 2+\varepsilon_{0}\,,\;|\lambda_{1}(\boldsymbol{Y})-\lambda_{*}|\leq\varepsilon_{0}\big{\}}\,.$ (81)
By the Bai-Yin law and Gaussian concentration (plus the above concentration of $\lambda_{1}$), $\mathbb{P}(\mathcal{E})\geq 1-2e^{-c(\varepsilon_{0})n}$ for some $c(\varepsilon_{0})>0$. Further, it is easy to check that $F(\boldsymbol{W})$ is Lipschitz on $\mathcal{E}$, whence the concentration of $\langle\boldsymbol{u},\boldsymbol{v}_{1}(\boldsymbol{W})\rangle^{2}$ follows by another application of Borell inequality.
Appendix K Proofs of reduction results
K.1 Proof of Theorem 5
The algorithm consists in running the discretized diffusion (4) with initialization $\hat{\boldsymbol{y}}_{t_{0}}=\boldsymbol{y}/\sigma^{2}$ at $t=t_{0}:=1/\sigma^{2}$. To avoid notational burden, we will assume $(T-t_{0})/\Delta$ to be an integer. Let $\hat{\boldsymbol{y}}_{t_{0}}^{*}$ be generated by the discretized diffusion with initialization at $\hat{\boldsymbol{y}}_{0}$ at $t=0$. Note that the distribution of $\hat{\boldsymbol{y}}_{t_{0}}$ is the same as the one of $t_{0}\boldsymbol{x}+\sqrt{t}\boldsymbol{g}$ and hence by Assumption $(b)$,
$W_{1}({\rm P}_{\hat{\boldsymbol{y}}_{t_{0}}},{\rm P}_{\hat{\boldsymbol{y}}^{*}_{t_{0}}})\leq W_{1}({\rm P}^{\hat{\boldsymbol{y}}}_{T,\Delta},{\rm P}^{\boldsymbol{y}}_{T})\leq\varepsilon\,.$ (82)
In other words there exists a coupling of $\hat{\boldsymbol{y}}^{*}_{t_{0}}$ and $\hat{\boldsymbol{y}}_{t_{0}}$ such that $\mathbb{E}\|\hat{\boldsymbol{y}}^{*}_{t_{0}}-\hat{\boldsymbol{y}}_{t_{0}}\|_{2}\leq\varepsilon$.
We extend this to a coupling of $(\hat{\boldsymbol{y}}^{*}_{t})_{t_{0}\leq t\leq T}$ and $(\hat{\boldsymbol{y}}_{t})_{t_{0}\leq t\leq T}$ in the obvious way: we generate the two trajectories according to the discretized diffusion (4) with the same randomness $\hat{\boldsymbol{z}}_{t}$. A simple recursive argument (using the Lipschitz property of $\hat{\boldsymbol{m}}$, in Assumption $(a)$) then yields
$\mathbb{E}\|\hat{\boldsymbol{y}}^{*}_{T}-\hat{\boldsymbol{y}}_{T}\|_{2}\leq\Big{(}1+L\Delta/d\Big{)}^{T/\Delta}\varepsilon\leq e^{LT/d}\varepsilon\,.$ (83)
(See for instance [MW23] or [AMS23] for examples of this calculation.) Let now $\boldsymbol{y}_{T}\stackrel{{\scriptstyle{\rm d}}}{{=}}T\boldsymbol{x}+\sqrt{T}\boldsymbol{g}^{\prime}$ for $(\boldsymbol{x},\boldsymbol{g}^{\prime})\sim\mu\otimes{\mathsf{N}}(\boldsymbol{0},\boldsymbol{I})$. Another application of Assumption $(a)$ implies that this can be coupled to $\hat{\boldsymbol{y}}^{*}_{T}$ so that $\mathbb{E}\|\boldsymbol{y}_{T}-\hat{\boldsymbol{y}}^{*}_{T}\|\leq\varepsilon$, and therefore
$\mathbb{E}\|\hat{\boldsymbol{y}}_{T}-\boldsymbol{y}_{T}\|_{2}\leq 2\,e^{LT/d}\varepsilon\,.$ (84)
As output, we return $\hat{\boldsymbol{x}}=\hat{\boldsymbol{y}}_{T}/T$. Using $\mathbb{E}\|\boldsymbol{y}_{T}-\boldsymbol{x}\|=\mathbb{E}\|\boldsymbol{g}\|/\sqrt{T}$ and $T=\theta d$,
$\mathbb{E}\|\boldsymbol{x}-\hat{\boldsymbol{x}}\|\leq 2e^{\theta L}\varepsilon+\frac{1}{\sqrt{\theta}}\,.$ (85)
Since the coupling has been constructed conditionally on $\boldsymbol{y}$, the claim (19) follows.
Finally, Eq. (17) follows by generating $N$ i.i.d. copies $\hat{\boldsymbol{x}}_{1},\dots,\hat{\boldsymbol{x}}_{N}$ using the above procedure, and letting $\hat{\boldsymbol{m}}(\boldsymbol{y})$ be their empirical average.